### v1

# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

Abstract
This paper investigates the failure of certain metric measure spaces to be infinitesimally Hilbertian or quasi-Riemannian manifolds, by constructing examples arising from a manifold $M$ endowed with a Riemannian metric $g$ that is possibly discontinuous, with $g,g^{-1}\in L^{\infty}_{\mathrm{loc}}$ and $g\in W^{1,p}_{\mathrm{loc}}$ for $p<\dim M-1$. footnotetext: MSC2020: 53C20, 53C23
Keywords. Riemannian manifolds of low regularity, metric measure spaces, infinitesimal Hilbertianity
Acknowledgements. The author gratefully acknowledges Andrea Mondino for his supervision and guidance, and thanks Alessandro Cucinotta for valuable discussions. The author is supported by a postgraduate scholarship from the Mathematical Institute at the University of Oxford. Further financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813, which funded the 2025 Trimester Program “Metric Analysis” in Bonn, is also acknowledged with gratitude.
1 Introduction
Manifolds with (semi)-Riemannian metrics of low regularity (in particular below $C^{2}$) have been of growing interest and importance in the field of geometric analysis. They may arise from gluing constructions, or in the case of Lorentzian signature (,+,,+)(-,+$\ldots$,+)( - , + , … , + ), serve to understand physically relevant phenomena from general relativity (see for instance [5]).
One approach to generalise Riemannian manifolds is to drop any underlying smoothness assumption and study metric measure spaces. A metric measure space is a triplet $(X,\mathsf{d},\mathfrak{m})$, where $(X,\mathsf{d})$ is a complete and separable metric space and $\mathfrak{m}$ is a $\sigma$-finite Borel measure on $X$ (playing the role of the volume measure). For a similar generalisation of Lorentzian manifolds see [22, 6].
Various concepts from Riemannian geometry can be generalised to metric measure spaces such as synthetic notions of the Ricci curvature bounded below by $K$ and the dimension bounded above by $N$ for some $K\in\mathbb{R}$ and $N\in[1,\infty]$; the curvature dimension condition $\mathsf{CD}(K,N)$ defined by Lott-Villani [25, 24] and Sturm [35, 36].
In this paper however, we will focus on the generalised first order calculus on metric measure spaces. It emerged through the search for a criterion on metric measure spaces that distinguishes those arising from Finslerian structures from metric measure spaces arising from Riemannian structures. This lead to the study of Sobolev calculus on metric measure spaces. After the approaches from Cheeger [7], Hajłasz [19], and Shanmungalingam [32], Ambrosio, Gigli and Savaré introduced a Sobolev calculus based on the notion of test plans in [2] that they proved to be equivalent to the approaches in [7] and [32], leading to the definition of the Sobolev space $W^{1,2}(X)$, and a Laplace operator.
In [14], Gigli studied conditions under which the Laplacian is linear and proposed the notion of infinitesimal Hilbertianity. A metric measure space is infinitesimally Hilbertian, if the associated Sobolev space $W^{1,2}(X)$ is a Hilbert space, or equivalently if the Cheeger energy (a generalisation of the Dirichlet energy to metric measure spaces) is a quadratic form. Then the Laplacian is automatically linear and the theory on Dirichlet forms provides various tools to better understand the Laplacian and the heat flow.
Examples of infinitesimally Hilbertian metric measure spaces include geodesically complete weighted Riemannian manifolds [26], locally $\mathrm{CAT}(\kappa)$-spaces equipped with a positive and locally finite Radon measure [11], and sub-Riemannian manifolds equipped with an arbitrary Radon measure [23]. Relevant notions to study the Cheeger energy, are the metric speed of absolutely continuous curves ([1], [21]) and the slope of Lipschitz continuous functions. In [3], Ambrosio, Gigli, and Savaré proposed the Riemannian curvature dimension condition, $\mathsf{RCD}(K,\infty)$, a refinement of the curvature dimension condition which combines the $\mathsf{CD}(K,\infty)$-condition with infinitesimal Hilbertianity. The $\mathsf{RCD}(K,N)$-condition for $N<\infty$ was proposed by Gigli in [13]. Using this terminology, an earlier result obtained by Ohta and Sturm in [30] states that $(\mathbb{R}^{n},\norm{\cdot},\mathcal{L}^{n})$ is an $\mathsf{RCD}(0,N)$-space if and only if $\norm{\cdot}$ arises from an inner product, hence in that case infinitesimal Hilbertianity indeed rules out Riemannian structures among Finslerian ones.
A smooth $d$-manifold $M$, equipped with a continuous Riemannian metric $g$ can naturally be seen as a metric measure space, where the measure is induced by the volume form (locally given via ${\rm d}\mathrm{vol}_{g}=\sqrt{\det g}\,{\rm d}\mathcal{L}^{d}$) and the Riemannian distance $\mathsf{d}_{g}$ is defined as the infimum over lengths of piecewise smooth curves (see (2.3)).
However, there are several relevant classes of Riemannian metrics that are non-continuous, for instance the Geroch-Traschen class of metrics [12], i.e. those Riemannian metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ and $g\in W^{1,2}_{\rm loc}(M)$. These regularity assumptions are minimal to compute the distributional Ricci curvature on $M$ (see [18], [16]) and have been studied in both Riemannian (see for example [17]) and Lorentzian signature (see [33]). Moreover, in the context of potential theory for uniformly elliptic operators, the even broader class of (Riemannian) metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ has been studied in [31], [34], [29] and [9].
If the metric tensor is only measurable, Norris (cf. [29]) and De Cecco-Palmieri (cf. [9]) defined a distance $\mathsf{d}_{g}$ that avoids the difficulty that piecewise smooth curves $\gamma$ might lie on a null set where $g$ is not well-defined (see Section 3).
This paper is driven by the following questions, which arise naturally from the results on smooth Riemannian manifolds:
Question 1 (see also [10]).
(Under which regularity assumptions on $g$) does $(M,\mathsf{d}_{g},g)$ satisfy that for any Lipschitz continuous function $f:M\to\mathbb{R}$ its slope equals the norm of its gradient almost everywhere, i.e. $|Df|=\limsup_{y\to x}\frac{|f(y)-f(x)|}{\mathsf{d}_{g}(x,y)}=|\nabla_{g}f|_{g}$ almost everywhere?
This question has been posed by De Giorgi in [10] as the positive answer would be a necessary condition for $(M,\mathsf{d}_{g},g)$ to be a quasi-Riemannian manifold, as defined in [10], see also Definition 3.6. A related question:
Question 2.
(Under which regularity assumptions on $g$) does the metric speed of “almost every” absolutely continuous curves equal the norm of its derivative, i.e. $\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})}{|h|}=|\dot{\gamma}_{t}|_{g}$ for almost every $t\in[0,1]$?
What exactly “almost every” curve refers to is clarified in Section 2.
Question 3.
(Under which regularity assumptions on $g$) is $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ infinitesimally Hilbertian?
This particular question refers back to the original purpose of infinitesimal Hilbertianity to identify the Riemannian-like structures among metric measure spaces.
It is already known that assuming $g\in C^{0}$ is sufficient for the second and third question to have positive answers (see [4], [28]). Moreover, it is known that for $g\in C^{0}$ and $f\in C^{1}$, the slope $|Df|$ of $f$ equals the norm of its gradient [28]. In this paper, we will provide examples showing the necessity of those regularity assumptions.
Outline of the paper and main results. In Section 2, we will recall some basic notions on the first order calculus on metric measure spaces and their explicit descriptions in the case of a manifold equipped with a continuous Riemannian metric. In Section 3, we recall the definition of the distance in the case of a non-continuous Riemannian metric by Norris and De Cecco-Palmieri, and briefly study absolutely continuous curves and test plans in that setting. In Section 4, we will construct Riemannian metrics $g$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not a quasi-Riemannian manifold and such that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian. We start with metrics that only satisfy $g,g^{-1}\in L^{\infty}_{\rm loc}$, providing a first example of a non-quasi Riemannian manifold in Subsection 4.1 and a first example of a non-infinitesimally Hilbertian space in Subsection 4.2. In Subsection 4.3, we refine the previous examples to a Sobolev metric $g$ such that $g,g^{-1}\in W^{1,p}_{\rm loc}\cap L^{\infty}_{\rm loc}$ for some $p<\dim M-1$. Questions 1 and 2 are then addressed in Subsection 4.3.1, where we prove:
Theorem 1.
For $d\geq 3$, $p\in[1,d-1)$ there exists a $d$-dimensional manifold $M$ and a Riemannian metric $g$ on $M$ such that $g,g^{-1}\in L^{\infty}$, $g\in W^{1,p}_{\rm loc}(M)$ and such that
  • (i)
    There exists a test plan that gives positive measure to curves for which the metric speed is strictly smaller than the norm of the derivative for almost every time $t\in[0,1]$.
  • (ii)
    There exists a function $f\in C^{1}(M)$ for which the minimal weak upper gradient and the metric slope (with respect to $\mathsf{d}_{g}$) are strictly larger than the norm of the gradient on a set of positive measure, i.e. $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
Question 3 is addressed in Subsection 4.3.2, where we prove:
Theorem 2.
For $d\geq 3$, $p\in[1,d-1)$ there exists a $d$-dimensional manifold $(M,g)$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ and $g\in W^{1,p}_{\rm loc}(M)$ such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
Notation
Given a manifold $M$, a Riemannian metric $g$, and a locally Lipschitz continuous function $f:M\to\mathbb{R}$, $|Df|$ will always denote its slope (with respect to the distance $\mathsf{d}_{g}$), $|Df|_{w}$ its minimal weak upper gradient, $\nabla_{g}f$ its gradient with respect to $g$ and $\nabla_{euc}f$ its differential with respect to chosen local coordinates (which will be specified).
Given an absolutely continuous curve $\gamma:[0,1]\to M$, we will for almost every $t\in[0,1]$ denote its metric speed at time $t$ by $|\dot{\gamma}_{t}|$, and its derivative at time $t$ by $\dot{\gamma}_{t}\in T_{\gamma_{t}}M$. To avoid confusion, if we take the norm of the derivative, we will always write $|\dot{\gamma}_{t}|_{g}$ for $\sqrt{g(\dot{\gamma}_{t},\dot{\gamma}_{t})}$ and $|\dot{\gamma}_{t}|_{euc}$ for $\sqrt{\langle\dot{\gamma}_{t},\dot{\gamma}_{t}\rangle_{euc}}$ (given specified local coordinates).
When working in $\mathbb{R}^{d}$, $e_{i}$ denotes the $i$-th unit vector, $\mathcal{L}^{d}$ the $d$-dimensional Lebesgue measure, and $Id$ denotes the identity matrix.
For any measure space $(X,\mathfrak{m})$ and $\mathfrak{m}$-measurable set $A\subset X$, $\mathbbm{1}_{A}$ denotes the characteristic function of $A$.
2 Basic definitions on metric measure spaces
In this section we will recall notions from first order calculus on metric measure spaces and their identification in metric measure spaces that arise from manifolds equipped with an at least continuous Riemannian metric.
2.1 First order calculus on metric measure spaces
We will briefly summarise some tools to define a generalised notion of modulus of gradients and Sobolev functions in metric measure spaces that was introduced by Cheeger in [7] and further analysed by Ambrosio, Gigli and Savaré [2]. The structure of this section has been inspired by [28].
Let $(X,\mathsf{d})$ be a complete and separable metric space. The slope (or local Lipschitz constant) of a real valued function $f:X\to\mathbb{R}$ is defined by
$|Df|(x):=\limsup_{y\to x}\frac{|f(x)-f(y)|}{\mathsf{d}(x,y)},$
if $\{x\}$ is not isolated, and 00 otherwise.
We endow $(X,\mathsf{d})$ with a non-negative $\sigma$-finite Borel measure $\mathfrak{m}$, obtaining the metric measure space $(X,\mathsf{d},\mathfrak{m})$. Throughout the rest of this work, we assume that there exists a bounded Borel Lipschitz map $V:X\to[0,\infty)$ such that (cf. [2, Ch. 4])
$\begin{split}V\ \mathrm{is\ bounded\ on\ each\ compact\ set\ }K\subset X\ \mathrm{and}\\ \int_{X}e^{-V^{2}}\,{\rm d}\mathfrak{m}\leq 1.\end{split}$ (2.1)
Definition 2.1.
Let $\gamma\in C([0,1],X)$ be a curve. Then we say that $\gamma$ is absolutely continuous if there exists a function $g\in L^{1}([0,1],[0,\infty))$ such that for all $0\leq s<t\leq 1$, it holds
$\mathsf{d}(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}g(r)\,{\rm d}r.$
There exists a minimal such function $g$, which we call the metric speed of $\gamma$ and denote it by $|\dot{\gamma}_{t}|$.
The metric speed of an absolutely continuous curve $\gamma:[0,1]\to X$ is given by $|\dot{\gamma}_{t}|:=\lim_{h\to 0}\frac{\mathsf{d}(\gamma_{t+h},\gamma_{t})}{|h|}$ which exists for almost every $t\in[0,1]$ [1], (see also [21]).
We next recall the notion of test plan and weak upper gradient. We will use the conventions of [15], after [2]. For any metric space $(Y,\mathsf{d}_{Y})$, we denote by $\mathcal{P}(Y)$ the set of Borel probability measures on $Y$. Given two topological spaces $X_{1},X_{2}$, a Borel-measurable map $T:X_{1}\to X_{2}$, and a Borel measure $\mu$ on $X_{1}$, the pushforward measure $T_{\#}\mu$ on $X_{2}$ is the Borel measure defined by $T_{\#}\mu(B)=\mu(T^{-1}(B))$ for any Borel set $B\subset X_{2}$.
Definition 2.2.
Let $\boldsymbol{\pi}\in\mathcal{P}(C([0,1],X))$. We say that $\boldsymbol{\pi}$ is a test plan if there exists a constant $C(\boldsymbol{\pi})>0$ such that
$(e_{t})_{\#}\boldsymbol{\pi}\leq C(\boldsymbol{\pi})\mathfrak{m}\ \mathrm{for\ all\ }t\in[0,1]$
and
$\int\int_{0}^{1}|\dot{\gamma}_{t}|^{2}\,{\rm d}t{\rm d}\boldsymbol{\pi}(\gamma)<\infty.$
We use the convention that if $\gamma$ is not absolutely continuous, then $\int_{0}^{1}|\dot{\gamma}_{t}|^{2}\,{\rm d}t=\infty$.
In Question 2, “almost every absolutely continuous curve” refers to a set of curves $\Gamma\subset C([0,1],M)$ such that for every test plan $\boldsymbol{\pi}$ as defined above, $\boldsymbol{\pi}(\Gamma)=1$.
Definition 2.3.
Given $f:X\to\mathbb{R}$ a $\mathfrak{m}$-measurable function, then a $\mathfrak{m}$-measurable function $G:X\to[0,\infty]$ is called a weak upper gradient of $f$, if
$\int|f(\gamma_{1})-f(\gamma_{0})|\,{\rm d}\boldsymbol{\pi}(\gamma)\,\leq\int\int_{0}^{1}G(\gamma_{t})|\dot{\gamma}_{t}|\,{\rm d}t{\rm d}\boldsymbol{\pi}(\gamma)<\infty,\quad\mathrm{for\ all\ test\ plans}\ \boldsymbol{\pi}.$ (2.2)
The discussion in [2], Prop. 5.9 and Def. 5.11 shows the existence of a weak upper gradient $|Df|_{w}$ such that $|Df|_{w}\leq G$ $\mathfrak{m}$-a.e. for all other weak upper gradients $G$. We will call it the minimal weak upper gradient of $f$.
Definition 2.4.
The Cheeger energy is defined in the class of $\mathfrak{m}$-measurable functions by
$\mathrm{Ch}(f):=\left\{\begin{array}[]{ll}\frac{1}{2}\int_{X}|Df|^{2}_{w}\,{\rm d}\mathfrak{m}&\mathrm{if}\ f\ \mathrm{has\ a\ weak\ upper\ gradient\ in}\ L^{2}(X,\mathfrak{m}),\\ \infty&\mathrm{otherwise},\end{array}\right.$
with proper domain $D(\mathrm{Ch})=\{f:X\to[0,\infty],\mathfrak{m}\mathrm{-measurable},\ \mathrm{Ch}(f)<\infty\}$.
Definition 2.5.
A metric measure space is called infinitesimally Hilbertian if $\mathrm{Ch}$ is a quadratic form.
2.2 Manifolds with continuous Riemannian metrics
Consider a $d$-dimensional manifold $M$ equipped with a Riemannian metric $g\in C^{0}(M)$. Then $g$ induces a volume measure whose density is locally given by ${\rm d}\mathrm{vol}_{g}=\sqrt{\det g}\,{\rm d}\mathcal{L}^{d}$. The distance induced by $g$ is given by
$\mathsf{d}_{g}(x,y)=\inf\Big{\{}\int_{0}^{1}|\dot{\gamma_{t}}|_{g}\,{\rm d}t:\,\gamma\ \mathrm{piecewise\ }C^{\infty},\gamma_{0}=x,\gamma_{1}=y\Big{\}}.$ (2.3)
In [4] the metric structure for such spaces $(M,\mathsf{d}_{g})$ has been thoroughly studied. In [28], the Cheeger energy and related notions were studied and expressed in terms of classical Sobolev spaces on manifolds (cf. [20]). We will recall some of these results to point out that in the case of a Riemannian manifold with a continuous metric, several notions from Subsection 2.1 can be identified with expected classical objects from calculus on manifolds.
Proposition 2.6 ([4] Prop. 4.1 and Thm 3.15).
The metric (2.3) turns $M$ into a length space.
A standard fact on the slope of functions is that if $f\in C^{1}(M)$, then $|Df|(x)=|\nabla_{g}f|_{g}(x)$ for every $x$.
Proposition 2.7 ([4], Proposition 4.10).
Let $\gamma:[0,1]\to M$ be an absolutely continuous curve. Then the metric speed $|\dot{\gamma_{t}}|$ coincides a.e. with $|\dot{\gamma}_{t}|_{g}=\sqrt{\langle\dot{\gamma}_{t},\dot{\gamma}_{t}\rangle_{g}}$, where $\dot{\gamma}_{t}$ denotes the (a.e. existing) derivative.
In the case of a manifold with a continuous Riemannian metric, it also turns out that the minimal weak upper gradient of a $C^{1}$-function $f$ can be identified as follows:
Proposition 2.8 ([28] Prop. 4.24).
Let $M$ be a smooth manifold, $g$ a $C^{0}$-Riemannian metric on $M$. Let $f$ be a $C^{1}$-function. Then
[…]
That was a crucial ingredient to conclude that:
Proposition 2.9 ([28], Cor. 4.27).
For a manifold $M$ equipped with a continuous Riemannian metric $g$, the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian.
3 The distance on manifolds with non-continuous Riemannian metrics
From now on we will consider a smooth $d$-dimensional manifold $M$ with a Riemannian metric $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}$. Then $(M,g)$ falls into the class of LIP-Riemannian manifolds (see for example [37, 9]). These manifolds and their metric structures, have been studied by Norris [29], and De Cecco and Palmieri [9], as well as Saloff-Coste [31] and Sturm [34] with emphasis on potential theory for uniformly elliptic operators. Note that LIP-manifolds shall not be confused with metric tensors whose coefficients are locally Lipschitz continuous.
If $g$ is only measurable, it might not be well-defined everywhere, hence the definition of the metric $\mathsf{d}_{g}$ from the continuous case (2.3), cannot be used for non-continuous Riemannian metrics. However, Norris [29] and De Cecco-Palmieri [9] gave a definition of a distance $\mathsf{d}_{g}$ on $M$ arising from $g$ that avoids this difficulty. We will briefly summarise their results and then study the metric speed of curves, test plans, and the Cheeger energy on the corresponding metric measure space.
In the case of a Riemannian manifold with $g,g^{-1}\in L^{\infty}$, Norris defines the distance $\mathsf{d}$ via
$\mathsf{d}(x,y):=\sup\left\{\left[w(x)-w(y)\right],\ w\mathrm{\ is\ Lipschitz\ continuous\ and\ }|\nabla_{g}w|_{g}\leq 1\ \mathrm{a.e.}\right\}.$
Moreover, for $\rho\in C_{c}^{\infty}(B_{1}^{\mathbb{R}^{d}}(0),[0,\infty))$, with $\int_{\mathbb{R}^{d}}\rho\,{\rm d}x=1$, define $\rho_{\varepsilon}:=\frac{1}{\varepsilon}\rho(\frac{x}{\varepsilon})$ and $g_{\varepsilon}:=\rho_{\varepsilon}*g$ (for details on the convolution on manifolds see [18, 16]). The metrics $g_{\varepsilon}$ are continuous and hence the distance $\mathsf{d}_{g_{\varepsilon}}$ is given as in (2.3). Norris then proves that for $\varepsilon\to 0$, $\mathsf{d}_{g_{\varepsilon}}$ converges to a distance $\mathsf{d}_{0}$ that is locally equivalent to the underlying Euclidean metric induced from a coordinate chart; in particular, [29, Theorem 3.6] states that $\mathsf{d}_{0}=\mathsf{d}=:\mathsf{d}_{g}$.
From the properties of $g$, it follows that all points outside a null set $N_{g}\subset M$ are Lebesgue points of $g,g^{-1}$. For any null set $N\subset M$, define the set
$\mathrm{Lip}_{N}(x,y,M):=\{\gamma\in C^{0,1}_{\rm{loc}}([0,1],M):\gamma(0)=x,\gamma(1)=y,\mathcal{L}^{1}(\{t\in[0,1]:\gamma(t)\in N\})=0\}.$ (3.1)
If $\gamma\in C([0,1],M)$ satisfies $\mathcal{L}^{1}(\{t\in[0,1]:\gamma(t)\in N\})=0$, we say that $\gamma$ is transversal to $N$ and write $\gamma\perp N$.
The reference measure is, as in the continuous case, given by ${\rm d}\mathrm{vol}_{g}=\sqrt{\det g}{\rm d}\mathcal{L}^{d}$. From now on, we will make a further assumption (see [9, (1.13)]):
Assumption 3.1.
The manifold $(M,g)$ satisfies one of the following:
  • (i)
    The manifold $M$ has finite $\mathrm{vol}_{g}$-volume.
  • (ii)
    We can cover $M$ with coordinate charts such that the smallest eigenvalue of $g$ is uniformly bounded below in these charts and $M$ is complete or the interior of a complete manifold (where the completeness is with respect to the distance induced by charts).
Finally, De Cecco and Palmieri proved in [9, Theorem (2.18) and Theorem (6.1)] that under Assumption 3.1, it holds $\mathsf{d}_{g}=\boldsymbol{\delta}$, where the distance $\boldsymbol{\delta}$ is given by
$\boldsymbol{\delta}(x,y)=\sup_{N\subset M,\mathrm{vol}_{g}(N)=0}\boldsymbol{\delta}_{N}(x,y),$ (3.2)
where
$\boldsymbol{\delta}_{N}:=\inf\{L_{g}(\gamma),\gamma\in\mathrm{Lip}_{N}(x,y,M)\},$ (3.3)
and $L_{g}(\gamma)=\int_{0}^{1}|\dot{\gamma}_{t}|_{g}{\rm d}t$. In [8, Theorem 4.4], it is shown that for each null set $N_{0}\subset M$, $\boldsymbol{\delta}=\boldsymbol{\delta}_{0}:=\sup_{N,\mathrm{vol}_{g}(N)=0}\boldsymbol{\delta}_{N\cup N_{0}}$.
Proposition 3.2.
Let $\gamma:[0,1]\to M$ be a $\mathsf{d}_{g}$-geodesic such that $\gamma\perp N_{g}$. Then, for almost every $t\in[0,1]$, it holds
$|\dot{\gamma}_{t}|\leq|\dot{\gamma}_{t}|_{g}.$
Proof.
For a Lipschitz curve $\gamma\perp N_{g}$, and any $t\in[0,1]$, $h\neq 0$, we have that
$\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})=\lim_{\varepsilon\to 0}\mathsf{d}_{g_{\varepsilon}}(\gamma_{t},\gamma_{t+h})\leq\lim_{\varepsilon\to 0}\int_{0}^{h}|\dot{\gamma}_{t+\tau}|_{g_{\varepsilon}}\,{\rm d}\tau=\int_{0}^{h}|\dot{\gamma}_{t+\tau}|_{g}\,{\rm d}\tau,$
where the last equality follows from the dominated convergence theorem and the fact that $\rho_{\varepsilon}*g\to g$ pointwise outside $N_{g}$. Now the characterisation $|\dot{\gamma}_{t}|=\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})}{|h|}$ a.e. together with the Lebesgue differentiation theorem yields the proposition. ∎
There are more observations that can be made on the metric speed of absolutely continuous curves. The first one is that the metric speed depends on the derivative of the curve.
Proposition 3.3.
Let $M$ be a smooth manifold and $g$ a Riemannian metric such that $g,g^{-1}\in L^{\infty}_{\rm loc}$. Let $\gamma:[0,1]\to M$ be a Lipschitz continuous curve and $t\in[0,1]$ such that the metric speed $|\dot{\gamma}_{t}|$ exists at $t$ and $\gamma$ is differentiable at $t$. Let $c:[0,1]\to M$ be another curve such that $c_{t}=\gamma_{t}$, $c$ is differentiable at $t$ and $\dot{c}_{t}=\dot{\gamma}_{t}$. Then the metric speed of $|\dot{c}_{t}|$ of $c$ at $t$ exists and $|\dot{c}_{t}|=|\dot{\gamma}_{t}|$.
Proof.
We work in local coordinates in $\mathbb{R}^{d}$ and write $\gamma_{t}=c_{t}=:p$ and $\dot{c}_{t}=\dot{\gamma}_{t}=v$. Furthermore, we may restrict to a small neighbourhood $U$ around $p$ and assume that there exist $\lambda,\Lambda>0$ such that $\lambda Id\leq g\leq\Lambda Id$. Then, by the definition of the derivative, we get that
$\gamma_{t+h}=p+hv+a_{\gamma}(h),\ c_{t+h}=p+hv+a_{c}(h)\in\mathbb{R}^{d},$
where
$\lim_{h\to 0}\frac{|a_{\gamma}(h)|_{euc}+|a_{c}(h)|_{euc}}{|h|}=0.$ (3.4)
Now, we can compute that
$\mathsf{d}_{g}(\gamma_{t+h},c_{t+h})\leq\sqrt{\Lambda}(|a_{\gamma}(h)|_{euc}+|a_{c}(h)|_{euc}),$ (3.5)
and hence
$0\displaystyle 0\leq\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t+h},c_{t+h})}{|h|}\leq\lim_{h\to 0}\sqrt{\Lambda}\frac{|a_{\gamma}(h)|_{euc}+|a_{c}(h)|_{euc}}{|h|}=$ 0 .
But then
$\limsup_{h\to 0}\bigg{|}\frac{\mathsf{d}_{g}(c_{t},c_{t+h})}{|h|}-\frac{\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})}{|h|}\bigg{|}\leq\limsup_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t+h},c_{t+h})}{|h|}=0.$
Next, we establish that the metric speed depends locally Lipschitz-continuously on the derivative of a curve.
Proposition 3.4.
Let $\gamma^{1},\gamma^{2}:[0,1]\to M$ be two Lipschitz continuous curves that are contained in a compact set $K$ inside one coordinate patch $V$ and $t\in[0,1]$ such that $\gamma^{1}_{t}=\gamma^{2}_{t}$ and such that the metric speeds $|\dot{\gamma}^{1}_{t}|$ and $|\dot{\gamma}^{2}_{t}|$ as well as their derivatives $\dot{\gamma}^{1}_{t},\dot{\gamma}^{2}_{t}\in T_{\gamma^{1}_{t}}M$ exist. Then there exists a constant $C=C(K)$ such that
$||\dot{\gamma}^{1}_{t}|-|\dot{\gamma}^{2}_{t}||\leq C|\dot{\gamma}^{1}_{t}-\dot{\gamma}^{2}_{t}|_{g}.$ (3.6)
Proof.
We may assume to be working in $\mathbb{R}^{d}$. Now there exist $\lambda,\Lambda>0$ such that $\lambda Id\leq g\leq\Lambda Id$ on $K$. Denote $v:=\dot{\gamma}^{1}_{t}$ and $w:=\dot{\gamma}^{2}_{t}$. By translation we may assume that $t=0$ and denote $p:=\gamma^{1}_{0}=\gamma^{2}_{0}$. By the previous proposition, we may assume that $\gamma^{1}_{t}=p+tv$ and $\gamma^{2}_{t}=p+tw$. But then
$\Big{|}\frac{\mathsf{d}_{g}(\gamma^{1}_{0},\gamma^{1}_{h})}{|h|}-\frac{\mathsf{d}_{g}(\gamma^{2}_{0},\gamma^{2}_{h})}{|h|}\Big{|}$ $\leq\frac{\mathsf{d}_{g}(p+hv,p+hw)}{|h|}\leq\sqrt{\Lambda}\frac{\mathsf{d}_{euc}(p+hv,p+hw)}{|h|}$
$=\sqrt{\Lambda}|v-w|_{euc}\leq\sqrt{\frac{\Lambda}{\lambda}}|v-w|_{g}=\sqrt{\frac{\Lambda}{\lambda}}|\dot{\gamma}^{1}_{t}-\dot{\gamma}^{2}_{t}|_{g}.$
Lemma 3.5.
Let $\boldsymbol{\pi}\in\mathcal{P}(C([0,1],M))$ be a test plan and $N\subset M$ a $\mathrm{vol}_{g}$-null set. Then, it holds that
$\boldsymbol{\pi}\Big{(}\{\gamma\in C([0,1],M):\gamma\perp N\}\Big{)}=1.$ (3.7)
Proof.
We argue by contradiction. Assume that $\boldsymbol{\pi}\Big{(}\{\gamma\in C([0,1],M):\gamma\perp N\}\Big{)}<1$. Then as
$\{\gamma\in C([0,1],M):\gamma\not\perp N\}=\bigcup_{n\in\mathbb{N}}\{\gamma\in C([0,1],M):\mathcal{L}^{1}(\gamma^{-1}(N))\geq\frac{1}{n}\},$
there exists an $n\in\mathbb{N}$ and an $\varepsilon>0$ such that
$\boldsymbol{\pi}(\{\gamma\in C([0,1],M):\mathcal{L}^{1}(\gamma^{-1}(N))\geq\frac{1}{n}\})=\varepsilon>0.$ (3.8)
Denote $\{\gamma\in C([0,1],M):\mathcal{L}^{1}(\gamma^{-1}(N))\geq\frac{1}{n}\}=:B_{n}$ and note that for each curve $\gamma\in B_{n}$, it holds
$\int_{0}^{1}\mathbbm{1}_{N}(\gamma_{t})\,{\rm d}t\geq\frac{1}{n},$ (3.9)
hence, with Fubini’s theorem, we get that
$\int_{0}^{1}\int_{C([0,1],M)}\mathbbm{1}_{N}(\gamma_{t})\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t=\int_{C([0,1],M)}\int_{0}^{1}\mathbbm{1}_{N}(\gamma_{t})\,{\rm d}t{\rm d}\boldsymbol{\pi}(\gamma)\geq\frac{\varepsilon}{n}>0.$ (3.10)
By the definition of a test plan, we have that for all $t\in[0,1]$ it holds $(e_{t})_{\#}\boldsymbol{\pi}\ll\mathrm{vol}_{g}$. As $N$ is a ${\rm d}\mathrm{vol}_{g}$-null set, we get that for all $t\in[0,1]$, it holds
$\int_{M}\mathbbm{1}_{N}(y)\,{\rm d}(e_{t})_{\#}\boldsymbol{\pi}(y)=0.$
That gives
$\int_{0}^{1}\int_{C([0,1],M)}\mathbbm{1}_{N}(\gamma_{t})\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t=\int_{0}^{1}\int_{M}\mathbbm{1}_{N}(y)\,{\rm d}(e_{t})_{\#}\boldsymbol{\pi}(y){\rm d}t=0,$
which contradicts (3.10). This proves the Lemma. ∎
We conclude this section with the definition of a quasi-Riemannian manifold proposed by De Giorgi in [10], that we will refer to later:
Definition 3.6.
Given a manifold $M$, a distance $\mathsf{d}$ on $M$, and a measurable positive definite inner product $g$ on $TM$, we call $(M,\mathsf{d},g)$ a quasi-Riemannian manifold if there exists an atlas $\mathcal{A}=\{(W,\psi)\}$ such that the following conditions hold:
  • (i)
    In each $(W,\psi)$, there exist constants $\lambda,\Lambda>0$ such that for any $x,y\in\psi(W)$, it holds
    $\lambda\leq\sum_{i,j}(\psi_{*}g)^{ij}\partial_{x^{i}}\mathsf{d}(\psi^{-1}(x),\psi^{-1}(y))\partial_{x^{j}}\mathsf{d}(\psi^{-1}(x),\psi^{-1}(y))\leq\Lambda.$
  • (ii)
    For any $f\in C^{0}_{c}(W)$, it holds
    $\int_{W}f{\rm d}\mathcal{H}^{d}_{\mathsf{d}}=\int_{W}f\sqrt{\det g}\,{\rm d}\mathcal{L}^{d},$
    where $\mathcal{H}^{d}_{\mathsf{d}}$ denotes the $d$-dimensional Hausdorff measure with respect to the distance $\mathsf{d}$.
  • (iii)
    For any Lipschitz continuous $f:M\to\mathbb{R}$, and almost every $x\in M$ it holds
    $\limsup_{y\to x}\frac{|f(x)-f(y)|}{\mathsf{d}(x,y)}=|\nabla_{g}f|_{g}.$
4 Construction of examples
In this section, we will construct Riemannian manifolds with non-continuous metrics that together with the distance $\mathsf{d}_{g}$ as defined in the previous section, do not give rise to quasi-Riemannian manifolds or infinitesimally Hilbertian metric measure spaces.
4.1 A first example of a non-quasi Riemannian manifold
We will now construct a metric $g$ on $\mathbb{R}^{d}$ such that $g,g^{-1}\in L^{\infty}$ and such that there exists a smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ such that $|Df|_{w},|Df|>|\nabla_{g}f|_{g}$ on a set of positive measure.
Let $q_{i}\in\mathbb{Q}$ be a counting of $\mathbb{Q}$. Pick a $\kappa\in(0,\frac{1}{8})$ and construct the following open set
$O:=\bigcup_{i=1}^{\infty}\Big{(}q_{i}-\frac{\kappa}{2^{i}},q_{i}+\frac{\kappa}{2^{i}}\Big{)}.$ (4.1)
Then $\mathcal{L}^{1}(O)\leq 2\kappa$ and $O$ is dense in $\mathbb{R}$. We now define
$\theta:=(2-\mathbbm{1}_{O})\in L^{\infty}(\mathbb{R}),$
$g(x_{1},\ldots,x_{d}):=\theta(x_{1})Id\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Then $(\mathbb{R}^{d},g)$ satisfies (ii) from Assumption 3.1 and $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies condition (2.1). We observe that
$(2-2\kappa)\leq\int_{(0,1)}\theta\,{\rm d}\mathcal{L}^{1}$ $\leq\frac{3}{2}\mathcal{L}^{1}\Big{(}\big{\{}\theta<\frac{3}{2}\big{\}}\cap(0,1)\Big{)}+2\mathcal{L}^{1}\Big{(}\big{\{}\theta\geq\frac{3}{2}\big{\}}\cap(0,1)\Big{)}$
$=\frac{3}{2}\Big{(}1-\mathcal{L}^{1}\Big{(}\big{\{}\theta\geq\frac{3}{2}\big{\}}\cap(0,1)\Big{)}\Big{)}+2\mathcal{L}^{1}\Big{(}\big{\{}\theta\geq\frac{3}{2}\big{\}}\cap(0,1)\Big{)}.$
It follows that
$\mathcal{L}^{1}\Big{(}\big{\{}\theta\geq\frac{3}{2}\big{\}}\cap(0,1)\Big{)}\geq 1-4\kappa\geq\frac{1}{2}.$ (4.2)
Now for any point $y\in(0,1)\times\mathbb{R}^{d-1}$ consider the curve $\gamma^{y}:[0,1]\to\mathbb{R}^{d},t\mapsto y+te_{2}$.
Claim. For all $t\in(0,1)$, $h\in(0,1-t)$ it holds
$\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})=h.$ (4.3)
Proof of the claim. Fix $t,h$ as above and fix an $\varepsilon>0$. Note that from $g\geq Id$, we get that $\mathsf{d}_{g}\geq\mathsf{d}_{euc}$, hence $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\geq h$. Thus, it remains to prove that $\mathsf{d}_{g}(\gamma^{y}_{t},\gamma^{y}_{t+h})\leq h$. There exists a $q\in\mathbb{Q}\cap(0,1)$ such that $|x_{1}(y)-q|\leq\varepsilon$. Let $i\in\mathbb{N}$ be such that $q_{i}=q$ in the chosen counting. Then define the curves
$\lambda_{1}:[0,1]\to\mathbb{R}^{d},s\mapsto\gamma^{y}_{t}+s(q-x_{1}(y))e_{1},$
$\lambda_{2}:[0,h]\to\mathbb{R}^{d},s\mapsto\gamma^{y}_{t}+(q-x_{1}(y))e_{1}+se_{2},$
$\lambda_{3}:[0,1]\to\mathbb{R}^{d},s\mapsto\gamma^{y}_{t}+(q-x_{1}(y))e_{1}+he_{2}-s(q-x_{1}(y))e_{1}.$
Denote by $\lambda$ the concatenation of those three curves and note that it yields a piecewise smooth curve that connects $\gamma^{y}_{t}$ and $\gamma^{y}_{t+h}$. Now, for all $\varsigma>0$, it holds $g_{\varsigma}=g*\rho_{\varsigma}\leq 2Id$, so we get that $L_{g_{\varsigma}}(\lambda_{1})+L_{g_{\varsigma}}(\lambda_{3})\leq 4\varepsilon$. For $\varsigma\leq\frac{\kappa}{2^{i+2}}$ that $g_{\varsigma}|_{\{x_{1}=q\}}\equiv Id$, hence $L_{g_{\varsigma}}(\lambda_{2})=h$. Thus,
$\lim_{\varsigma\to 0}\mathsf{d}_{g_{\varsigma}}(\gamma^{y}_{t},\gamma^{y}_{t+h})\leq\lim_{\varsigma\to 0}L_{g_{\varsigma}}(\lambda)\leq h+4\varepsilon.$
As $\varepsilon$ was arbitrary, this proves the claim.
By (4.2), we can choose a set $E^{\prime}\subset\{\theta\geq\frac{3}{2}\big{\}}\cap(0,1)$ such that $\mathcal{L}^{1}(E^{\prime})>0$. Denote $E:=E^{\prime}\times(0,1)^{d-1}$. Define the test plan $\boldsymbol{\pi}\in\mathcal{P}(C^{\infty}((0,1),\mathbb{R}^{d}))$ via
${\rm d}\boldsymbol{\pi}(\gamma):=\left\{\begin{array}[]{ll}\frac{1}{\mathcal{L}^{d}(E)}{\rm d}\mathcal{L}^{d}(y)&\mathrm{if}\ \gamma=\gamma^{y},\ \mathrm{for\ some}\ y\in E,\\ 0&\mathrm{otherwise}.\end{array}\right.$
Note that this is indeed a feasible test plan, as ${\rm d}\mathcal{L}^{d}\leq{\rm d}\mathrm{vol}_{g}\leq 2^{\frac{d}{2}}{\rm d}\mathcal{L}^{d}$. But now for all $y\in E$, almost all $t\in[0,1]$, we get that $g(\gamma^{y}_{t})\geq\frac{3}{2}Id$ and $\dot{\gamma}^{y}_{t}=e_{2}$ and $|\dot{\gamma}^{y}_{t}|_{g}\geq\sqrt{\frac{3}{2}}$. But by (4.3), we get that the metric speed $|\dot{\gamma}^{y}_{t}|=1$ for all $y,t$, hence we found a test plan that gives positive measure to curves whose metric speed and the norm of the derivative are not equal.
We will now give an example of a function $f\in C^{\infty}(\mathbb{R}^{d})$ for which the norm of the gradient is strictly smaller than the minimal weak upper gradient and the metric slope on a set of points with positive measure. The function is given by $f:\mathbb{R}^{d}\to\mathbb{R},x=(x_{1},\ldots,x_{d})\mapsto x_{2}$. Then for all $y\in E$, we have that $|\nabla_{g}f|_{g}(y)\leq\sqrt{\frac{2}{3}}$. But now for any point $y\in\mathbb{R}^{d}$, we have that
$|Df|(y)=\limsup_{z\to y}\frac{|f(z)-f(y)|}{\mathsf{d}_{g}(z,y)}\geq\lim_{h\to 0}\frac{|f(y+he_{2})-f(y)|}{\mathsf{d}_{g}(y+he_{2},y)}=\lim_{h\to 0}\frac{|h|}{|h|}=1>\sqrt{\frac{2}{3}}.$ (4.4)
This shows that for $y\in E^{\prime}\times\mathbb{R}^{d-1}$, it holds $|Df|(y)>|\nabla_{g}f|_{g}(y)$.
Finally, the test plan $\boldsymbol{\pi}$ shows, that $|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive measure. Indeed, we note that
$\int_{C([0,1],\mathbb{R}^{d})}|f(\gamma_{1})-f(\gamma_{0})|\,{\rm d}\boldsymbol{\pi}(\gamma)=\int_{E}\frac{1}{\mathcal{L}^{d}(E)}\,{\rm d}\mathcal{L}^{d}=1.$
On the other hand, we have that
$\int_{C([0,1],\mathbb{R}^{d})}\int_{0}^{1}|\nabla_{g}f(\gamma_{t})|_{g}|\dot{\gamma}_{t}|\,{\rm d}t{\rm d}\boldsymbol{\pi}(\gamma)\leq\int_{E}\int_{0}^{1}\sqrt{\frac{2}{3}}\frac{1}{\mathcal{L}^{d}(E)}\,{\rm d}t{\rm d}\mathcal{L}^{d}=\sqrt{\frac{2}{3}}.$
Hence, we have proved the following:
Proposition 4.1.
For $d\geq 2$, there exists a Riemannian metric $g$ on $\mathbb{R}^{d}$ such that $g,g^{-1}\in L^{\infty}$ and such that
  • (i)
    There exists a test plan that gives positive measure to curves for which the metric speed is strictly smaller than the norm of the derivative for almost every time $t\in[0,1]$.
  • (ii)
    There exists a function $f\in C^{1}(\mathbb{R}^{d})$ for which the minimal weak upper gradient and the metric slope (with respect to $\mathsf{d}_{g}$) are strictly larger than the norm of the gradient on a set of positive measure, i.e. $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
4.2 A first example of non-infinitesimal Hilbertianity
In this subsection, we will more closely analyse the infinitesimal structure of a manifold similar to the previous example. This following example metric resembles the previous one, with the sole difference, that we only perturb the metric inside the interior of the unit cube $(0,1)^{d}\subset\mathbb{R}^{d}$ and leave it constant outside that set. We will prove that the resulting metric measure space is not infinitesimally Hilbertian.
Let $q_{i}\in\mathbb{Q}$ be a counting of $\mathbb{Q}\cap(0,1)$, $\kappa\in(0,\frac{1}{8})$ and define
$O:=\bigcup_{i=1}^{\infty}\Big{(}q_{i}-\frac{\kappa}{2^{i}},q_{i}+\frac{\kappa}{2^{i}}\Big{)}.$ (4.5)
Then $\mathcal{L}^{1}(O)\leq 2\kappa$ and $O$ is dense in $(0,1)$. Define
$\theta:=(2-\mathbbm{1}_{O})\in L^{\infty}((0,1)),$
$g:=2\mathbbm{1}_{((0,1)^{d})^{c}}Id+\mathbbm{1}_{(0,1)^{d}}\theta(x_{1})Id\in L^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d\times d}).$
Note that $(\mathbb{R}^{d},g)$ satisfies (ii) from Assumption 3.1, $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1), and $g$ is constant outside the precompact set $(0,1)^{d}$. Inside $(0,1)^{d}$, $g$ the metric space arising from $(M,g)$ will behave just as in the previous subsection, we only restrict this behaviour to a compact set to conclude the lack of infinitesimal Hilbertianity.
To see that we will first quantitatively determine the difference between the metric speed of curves and the norms of their derivatives.
We start with a preparatory lemma:
Lemma 4.2.
Let $f:[0,1]\to\mathbb{R}$ be a Lipschitz continuous function and $a<b\in\mathbb{R}$. Then
$\int_{f^{-1}((a,b))}f^{\prime}\,{\rm d}t\in[a-b,b-a].$ (4.6)
Proof.
We define the function $h:[0,1]\to\mathbb{R}$ as
$h(t):=\max(a,\min(f(t),b)).$
Then $h$ is Lipschitz continuous and admits a weak derivative $h^{\prime}$. Moreover, in the open set $f^{-1}((a,b))\subset[0,1]$, we have that $f=h$, hence $h^{\prime}=f^{\prime}$ almost everywhere in $f^{-1}((a,b))$. Now, on $f^{-1}((-\infty,a])$ we have that $h\equiv a$.
Claim. For almost every $t\in f^{-1}((-\infty,a])$, it holds $h^{\prime}=0$.
Proof of the claim. Indeed, choose a Lebesgue point $t$ of both $f^{-1}((-\infty,a])$ and $h^{\prime}$ and suppose for a contradiction that $h^{\prime}(t)\neq 0$. Then there exists a neighbourhood $U\ni t$ such that $h(s)\neq h(t)$ for every $s\in U\setminus\{t\}$. But as $t$ is a Lebesgue point of $f^{-1}((-\infty,a])$, we have that $(f^{-1}((-\infty,a])\cap U)\setminus\{t\}\neq\emptyset$. Hence, there exists a point in $s\in f^{-1}((-\infty,a])$ such that $h(s)\neq h(t)$, which contradicts the definition of $h$. As those Lebesgue points have full measure, this proves the claim.
We similarly get that on $f^{-1}([b,\infty))$, it holds $h^{\prime}\equiv 0$. Then
$\int_{f^{-1}((a,b))}f^{\prime}\,{\rm d}t=\int_{0}^{1}h^{\prime}\,{\rm d}t=h(1)-h(0).$ (4.7)
The lemma follows as $h(0),h(1)\in[a,b]$. ∎
By Proposition 3.3, we get that for any curve $\gamma$ and $t\in[0,1]$ such that $\gamma_{t}\in O\times(0,1)^{d-1}\cup((0,1)^{d})^{c}$ and such that $\gamma$ is differentiable at $t$, it holds $|\dot{\gamma}_{t}|=|\dot{\gamma}_{t}|_{g(\gamma_{t})}$.
Next suppose that $y\in(0,1)^{d}$ is a Lebesgue point of $\{g=2\}$. In this argument, we will work with the definition of the metric $\mathsf{d}_{g}$ via (3.2). We will at first determine the shape and length of an almost distance minimising curve between two points that are close to each other. Fix an $\varepsilon>0$. Then, by the Lebesgue differentiation theorem, there exists a $\delta>0$ such that for each open interval $I\subset\mathbb{R}$ with $x_{1}(y)\in I$, and $\mathcal{L}^{1}(I)\leq\delta$, it holds
$\frac{1}{\mathcal{L}^{1}(I)}\int_{I}|\theta-2|\,{\rm d}\mathcal{L}^{1}\leq\varepsilon.$ (4.8)
Fix the null set $N^{(1)}_{g}\subset\mathbb{R}$ as the set of non-Lebesgue points of $\theta$ and define $N_{g}:=N^{(1)}_{g}\times\mathbb{R}^{d-1}$. We may assume that $O\cap N^{(1)}_{g}=\emptyset$.
Let $z\in(0,1)^{d}$ be such that $\mathsf{d}_{euc}(y,z)\leq\frac{1}{8}\mathrm{dist}_{euc}(y,\partial(0,1)^{d})$. Then any $2$-shortest curve connecting them, i.e. any curve $\gamma:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ such that $\gamma\perp N_{g}$ and $L_{g}(\gamma)\leq 2\mathsf{d}_{g}(y,z)$, has to lie inside $(0,1)^{d}$. Let $\gamma:[0,1]\to(0,1)^{d}$ be a Lipschitz continuous curve transversal to $N_{g}$ such that $\gamma(0)=y$ and $\gamma(1)=z$. We will now estimate $\mathsf{d}_{g}(y,z)$ depending on $z-y\in\mathbb{R}^{d}$. We assume that $x_{1}(z)\geq x_{1}(y)$ and $|x_{1}(z)-x_{1}(y)|<\delta$. The case $x_{1}(z)\leq x_{1}(y)$ is analogous.
Step 1: Reduction to curves with uniformly bounded $x_{1}$ coordinates.
Pick $a_{y}\leq x_{1}(y)$ and $a_{z}\geq x_{1}(z)$ such that $a_{y},a_{z}\in O\cap(0,1)$, and
$a_{z}-a_{y}\leq\delta,\ x_{1}(y)-a_{y}\leq\varepsilon,\ a_{z}-x_{1}(z)\leq\varepsilon,a_{z}-a_{y}\leq 2|z-y|_{euc}.$ (4.9)
Now define $\tilde{\gamma}:[0,1]\to\mathbb{R}^{d}$ via
$x_{1}(\tilde{\gamma})(t)=\max(a_{y},\min(a_{z},x_{1}(\gamma_{t}))),$
$x_{i}(\tilde{\gamma})(t)=x_{i}({\gamma}_{t})\ \mathrm{for}\ i=2,\ldots,d.$ (4.10)
Then $\tilde{\gamma}$ is Lipschitz continuous and transversal to $N_{g}$. Indeed, on $(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))$, we get that $\tilde{\gamma}\in\{a_{y},a_{z}\}\times(0,1)^{d-1}\subset N^{c}_{g}\cap\{g\equiv Id\}$ and on $(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))$, we have that $\gamma=\tilde{\gamma}$. Note that the methods from the proof of Lemma 4.2 yield that
$|\dot{\tilde{\gamma}}_{t}|_{euc}\leq|\dot{\gamma}_{t}|_{euc}\ \mathrm{a.e.\ on}\ [0,1].$ (4.11)
Now, we can compute
$L_{g}(\gamma)$ $=\int_{[0,1]}\sqrt{\theta(x_{1}(\gamma_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t$
$=\int_{(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))}\sqrt{\theta(x_{1}(\gamma_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t+\int_{(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))}\sqrt{\theta(x_{1}(\gamma_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t$
$\geq\int_{(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))}\sqrt{\theta(x_{1}(\gamma_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t+\int_{(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))}|\dot{\tilde{\gamma}}_{t}|_{euc}\,{\rm d}t$
$=\int_{(x_{1}\circ\gamma)^{-1}((a_{y},a_{z}))}\sqrt{\theta(x_{1}(\tilde{\gamma}_{t}))}|\dot{\tilde{\gamma}}_{t}|_{euc}\,{\rm d}t+\int_{(x_{1}\circ\gamma)^{-1}((-\infty,a_{y}]\cup[a_{z},\infty))}\sqrt{\theta(x_{1}(\tilde{\gamma}_{t}))}|\dot{\tilde{\gamma}}_{t}|_{euc}\,{\rm d}t$
$=\int_{[0,1]}\sqrt{\theta(x_{1}(\tilde{\gamma}_{t}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t=L_{g}(\tilde{\gamma}).$ (4.12)
Hence, we found a curve $\tilde{\gamma}$ with the same endpoints as $\gamma$ that is transversal to $N_{g}$ with bounded $x_{1}$ coordinates that is at most as long as $\gamma$. From now on, we write $\gamma:=\tilde{\gamma}$ with a slight abuse of notation.
Step 2. Reduction to curves consisting of two straight lines.
Define
$v_{(1)}:=\int_{(\theta\circ x_{1}\circ\gamma)^{-1}(1)}\dot{\gamma}_{t}\,{\rm d}t\in\mathbb{R}^{d},$
$v_{(2)}:=\int_{(\theta\circ x_{1}\circ\gamma)^{-1}(2)}\dot{\gamma}_{t}\,{\rm d}t\in\mathbb{R}^{d}.$
Note that $v_{(1)}+v_{(2)}={z}-y$ and by classical results about the Euclidean space it holds
$L_{g}({\gamma})$ $=\int_{[0,1]}\sqrt{\theta(x_{1}(\tilde{\gamma}))}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t$
$=\int_{(\theta\circ x_{1}\circ\gamma)^{-1}(2)}\sqrt{2}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t+\int_{(\theta\circ x_{1}\circ\gamma)^{-1}(1)}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t$
$\geq\sqrt{2}|v_{(2)}|_{euc}+|v_{(1)}|_{euc}.$ (4.13)
Write $O\cap(a_{y},a_{z}):=\bigcup_{n\in\mathbb{N}}I_{n}$, where $I_{n}$ are pairwise disjoint open intervals. By our choice of $\delta$ (see (4.8) and (4.9)), we get that $\mathcal{L}^{1}(O\cap(a_{y},a_{z}))\leq\varepsilon(a_{z}-a_{y})$. Now, by Lemma 4.2, we get the following estimate on the first component $(v_{(1)})_{1}$ of $v_{(1)}=\sum_{i=1}^{d}(v_{(1)})_{i}e_{i}$.
$|(v_{(1)})_{1}|=\Big{|}\sum_{n}\int_{(x_{1}\circ\gamma)^{-1}(I_{n})}\dot{(x_{1}\circ\gamma)}(t){\rm d}t\big{|}\leq\sum_{n}\mathcal{L}^{1}(I_{n})\leq\varepsilon|a_{z}-a_{y}|\leq 2\varepsilon|z-y|_{euc}.$
Now define
$w_{(1)}:=v_{(1)}-(v_{(1)})_{1}e_{1}\in\{0\}\times\mathbb{R}^{d-1},\ w_{(2)}=v_{(2)}+(v_{(1)})_{1}e_{1}\in\mathbb{R}^{d}.$
Then $w_{(1)}+w_{(2)}=z-y$ and
$|w_{(1)}-v_{(1)}|_{euc},|w_{(2)}-v_{(2)}|_{euc}\leq 2\varepsilon|z-y|_{euc}.$
Hence
$L_{g}(\tilde{\gamma})\geq\sqrt{2}|w_{(2)}|_{euc}+|w_{(1)}|_{euc}-6\varepsilon|z-y|_{euc},$ (4.14)
where $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{2},\ldots e_{d})$. By potentially applying a linear transform from $\mathrm{SO}(d-1)$, we may assume that $z-y\in\mathrm{span}(e_{1},e_{2})$. We reduce the right hand side of (4.14), if we orthogonally project both $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{2})$, so we may assume that $w_{(1)},w_{(2)}\in\mathrm{span}(e_{1},e_{2})$. Now, if $z-y=ae_{1}$, for some $a\in\mathbb{R}$ we get that $w_{(1)}=0$ and hence
$L_{g}(\tilde{\gamma})\geq\sqrt{2}a-6a\varepsilon.$ (4.15)
Otherwise, we may write $z-y=a(be_{1}+e_{2})$, where $a,b\in\mathbb{R}$, $a\neq 0$. Write $w_{(1)}=a(1-c)e_{2}$ and $w_{(2)}=a(be_{1}+ce_{2})$. Now, we want to choose the optimal parameter $c\in[0,1]$ such that
$\sqrt{2}|w_{(2)}|_{euc}+|w_{(1)}|_{euc},$
that is (4.14) without the error term, becomes minimal under the constraint that $w_{(1)}\in\mathrm{span}(e_{2})$. This reduces to finding $c$ such that
$f_{b}(c):=|1-c|+\sqrt{2}\sqrt{b^{2}+c^{2}}$ (4.16)
is minimised. From the definition of the function one can see that for $c\leq 0$ it holds $f_{b}(c)\geq f_{b}(0)$. Similarly, if $c\geq 1$, we get that $f_{b}(c)\geq f_{b}(1)$. Hence we only need to minimise for $c\in(0,1)$, in which $f_{b}$ is continuously differentiable and check the boundary values. We have that
$f_{b}(0)=1+\sqrt{2}|b|,f_{b}(1)=\sqrt{2(1+b^{2})}.$
Now the derivative for $c\in(0,1)$ is given by
$\frac{{\rm d}}{{\rm d}c}f_{b}(c)=-1+\frac{\sqrt{2}c}{\sqrt{b^{2}+c^{2}}}.$
The derivative can only vanish if $\sqrt{2}c_{*}=\sqrt{b^{2}+c_{*}^{2}}\iff c_{*}=|b|$. Now, if $|b|\leq 1$, we get that $c_{*}\in[0,1]$ and
$f_{b}(c_{*})=1-|b|+\sqrt{2(b^{2}+b^{2})}=1+|b|.$
Then, we get that $f_{b}(0),f_{b}(1)\geq f_{b}(c_{*})$. If $|b|\geq 1$, we only need to check the boundary values and get that $f_{b}(0)\geq f_{b}(1)$. Hence, for $|b|\geq 1$ the optimal value is $\sqrt{2(b^{2}+1)}$ and for $|b|\leq 1$, the optimal value is $1+|b|$.
More generally, writing $z-y=u=(u_{1},u^{\prime})$ where $u_{1}\in x_{1}(y)+(-\delta,\delta)\subset\mathbb{R}$ and $u^{\prime}\in\mathbb{R}^{d-1}$, we have that
$L_{g}(\gamma)\geq\left\{\begin{array}[]{ll}|u_{1}|+|u^{\prime}|_{euc}-6\varepsilon|u|_{euc}&\mathrm{if}\ |u_{1}|\leq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-6\varepsilon|u|_{euc}&\mathrm{otherwise}.\end{array}\right.$
More precisely, using that $\mathsf{d}_{g}\geq\boldsymbol{\delta}_{N_{g}}$, we get that for $u$ as above, it holds
$\mathsf{d}_{g}(y,y+u)\geq\left\{\begin{array}[]{ll}|u_{1}|+|u^{\prime}|_{euc}-6\sigma(p,u_{1})|u|_{euc}&\mathrm{if}\ |u_{1}|\leq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-6\sigma(p,u_{1})|u|_{euc}&\mathrm{otherwise},\end{array}\right.$ (4.19)
where
$\sigma(y,u_{1}):=\inf\{\varepsilon>0,\Big{|}\frac{1}{|I|}\int_{I}\theta\,{\rm d}x-2\Big{|}\leq\varepsilon,\forall I\subset\mathbb{R}\ \mathrm{open},x_{1}(y)\in I,\mathcal{L}^{1}(I)<|u_{1}|\}.$ (4.20)
As $x_{1}(y)$ is chosen as a Lebesgue point of $\theta$, we get that $\sigma(y,u_{1})\to 0$ as $|u_{1}|\to 0$.
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ and work with the regularisation. Fix $\epsilon>0$ For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2Id$. Hence, for all $\varsigma>0$, we have that
$\mathsf{d}_{g_{\varsigma}}(y,z)\leq\sqrt{2}|u|_{euc}.$
Moreover, we can find a rational $q\in\mathbb{Q}\cap(0,1)$ such that $|q-x_{1}(y)|\leq\epsilon|u|_{euc}$. Fix a $10<\alpha<\beta<$ 1 such that $U:=(\alpha,\beta)^{d-1}$ satisfies that $y,z\in(0,1)\times U$. Now for some $\varsigma_{\epsilon}>0$ small enough, we have that for $\varsigma\in(0,\varsigma_{\epsilon})$, it holds $g_{\varsigma}=Id$ on $\{q\}\times U$. If $|u_{1}|<|u^{\prime}|_{euc}$, write $b=\frac{u_{1}}{|u^{\prime}|_{euc}}$. Then define the curves
$\gamma_{1}:[0,1]\to\mathbb{R}^{d},t\mapsto y+t(q-x_{1}(y))e_{1},$
$\gamma_{2}:[0,1]\to\mathbb{R}^{d},t\mapsto y+(q-x_{1}(y))e_{1}+t(1-|b|)u^{\prime},$
$\gamma_{3}:[0,1]\to\mathbb{R}^{d},t\mapsto y+(q-x_{1}(y))e_{1}+(1-|b|)u^{\prime}-t(q-x_{1}(y))e_{1},$
$\gamma_{4}:[0,1]\to\mathbb{R}^{d},t\mapsto y+(1-|b|)u^{\prime}+t|b|u^{\prime}+tu_{1}e_{1}.$
We then get that
$L_{g_{\varsigma}}(\gamma_{1}),L_{g_{\varsigma}}(\gamma_{3})\leq\sqrt{2}|q-x_{1}(y)|\leq\sqrt{2}\epsilon|u|_{euc}.$
Moreover as $g_{\varsigma}\equiv Id$ on $\{q\}\times U\supset\gamma_{2}([0,1])$, we get that
$L_{g_{\varsigma}}(\gamma_{2})=(1-|b|)|u^{\prime}|_{euc}.$
Finally, we get that
$L_{g_{\varsigma}}(\gamma_{4})\leq\sqrt{2}\big{|}bu^{\prime}+u_{1}e_{1}\big{|}_{euc}=2|u_{1}|.$
Hence, for all $\varsigma$ small enough, the concatenation $\hat{\gamma}$ of the four curves satisfies
$\mathsf{d}_{g_{\varsigma}}(y,z)\leq L_{g_{\varsigma}}(\hat{\gamma})\leq 3\epsilon|u|_{euc}+\frac{|u^{\prime}|_{euc}-|u_{1}|}{|u^{\prime}|_{euc}}|u^{\prime}|_{euc}+2|u_{1}|=|u_{1}|+|u^{\prime}|_{euc}+3\epsilon|u|_{euc}.$ (4.21)
Hence, using $\mathsf{d}_{g}=\lim_{\varsigma\to 0}\mathsf{d}_{g_{\varsigma}}$ and the fact that $\epsilon$ was independent of $|u_{1}|$, sending $\epsilon\to 0$ yields that
$\mathsf{d}_{g}(y,y+u)\leq\alpha(u):=\left\{\begin{array}[]{ll}|u_{1}|+|u^{\prime}|_{euc}&\mathrm{if}\ |u_{1}|\leq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}&\mathrm{otherwise}.\end{array}\right.$ (4.24)
Now considering the curve $\gamma^{y,u}:[0,1]\to\mathbb{R}^{d},t\mapsto y+tu$, for $u\in\mathbb{R}^{d}$, (4.24) and (4.19) give that
$0\displaystyle 0\leq\lim_{h\to 0}\frac{\alpha(hu)-\mathsf{d}_{g}(\gamma^{y,u}_{0},\gamma^{y,u}_{h})}{|h|}\leq\lim_{h\to 0}\frac{\sigma(y,hu_{1})|hu|_{euc}}{|h|}=\lim_{h\to 0}\sigma(y,hu_{1})|u|_{euc}=$ 0 .
Hence
$\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma^{y,u}_{0},\gamma^{y,u}_{h})}{|h|}=\lim_{h\to 0}\frac{\alpha(hu)}{|h|}=\alpha(u).$
Now note that for $\mathcal{L}^{1}$-almost all $x\in\mathbb{R}$, it holds that $x\in O$ or $x$ is a Lebesgue point of $\{\theta=2\}$. Write $N^{(2)}\subset\mathbb{R}$ for the null set of points satisfy neither of that and define the $\mathcal{L}^{d}$-null set $N_{2}:=N^{(2)}\times\mathbb{R}^{d-1}\cup(\partial(0,1)^{d})$. We now have characterised the behaviour of the metric speed for almost all points in the manifold $M$.
Consider the smooth functions $f_{1},f_{2},f_{3},f_{4}:M=\mathbb{R}^{d}\to\mathbb{R}$ given by
$f_{1}:x\mapsto x_{1},f_{2}:x\mapsto x_{2},f_{3}:=f_{1}+f_{2},\ \mathrm{and}\ f_{4}:=f_{1}-f_{2}.$
Using the results on smooth Riemannian manifolds, we get that in $O\times(0,1)^{d-1}\cup([0,1]^{d})^{c}$ the minimal weak upper gradient of all functions equal the Euclidean norm of their gradients (see Proposition 2.8).
Now again pick a point $y\in(0,1)^{d}$ such that $x_{1}(y)$ is a Lebesgue point of $\{\theta=2\}$. Pick a vector $0\neq u\in\mathbb{R}^{d}$ and compute
$\partial_{u}f_{1}=u_{1},\partial_{u}f_{2}=u_{2},\partial_{u}f_{3}=u_{1}+u_{2},\partial_{u}f_{4}=u_{1}-u_{2}.$
Then
[…]
Now if $|u_{1}|\geq|u^{\prime}|_{euc}$, we have that $\alpha(u)=\sqrt{2}|u|_{euc}$, hence
$\frac{|\partial_{u}f_{1}|}{\alpha(u)}=\frac{|u_{1}|}{\sqrt{2}|u|_{euc}}\leq\frac{|u_{1}|}{\sqrt{2}|u_{1}|}=\frac{1}{\sqrt{2}},$ (4.25)
$\frac{|\partial_{u}f_{2}|}{\alpha(u)}=\frac{|u_{2}|}{\sqrt{2}|u|_{euc}}\leq\frac{|u_{2}|}{\sqrt{2}|u_{2}|}=\frac{1}{\sqrt{2}},$ (4.26)
$\frac{|\partial_{u}f_{3}|}{\alpha(u)}=\frac{|u_{1}+u_{2}|}{\sqrt{2}|u|_{euc}}\leq\frac{|u_{1}|+|u_{2}|}{\sqrt{2(u_{1}^{2}+u_{2}^{2})}}\leq 1,$ (4.27)
$\frac{|\partial_{u}f_{4}|}{\alpha(u)}=\frac{|u_{1}-u_{2}|}{\sqrt{2}|u|_{euc}}\leq\frac{|u_{1}|+|u_{2}|}{\sqrt{2(u_{1}^{2}+u_{2}^{2})}}\leq 1,$ (4.28)
Note that (4.27) and (4.28) follow from the fact that the quadratic mean is greater than or equal to the arithmetic mean. Moreover, we note that the equality in (4.25) can be achieved by choosing $u=u_{(1)}:=e_{1}$.
If $|u_{1}|<|u^{\prime}|_{euc}$, we have that $\alpha(u)=|u_{1}|+|u^{\prime}|_{euc}$, hence
$\frac{|\partial_{u}f_{1}|}{\alpha(u)}=\frac{|u_{1}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{1}|}{2|u_{1}|+(|u^{\prime}|_{euc}-|u_{1}|)}=\frac{1}{2},$ (4.29)
$\frac{|\partial_{u}f_{2}|}{\alpha(u)}=\frac{|u_{2}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{2}|}{|u_{2}|}=1,$ (4.30)
$\frac{|\partial_{u}f_{3}|}{\alpha(u)}=\frac{|u_{1}+u_{2}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{1}|+|u_{2}|}{|u_{1}|+|u_{2}|}=1,$ (4.31)
$\frac{|\partial_{u}f_{4}|}{\alpha(u)}=\frac{|u_{1}-u_{2}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{1}|+|u_{2}|}{|u_{1}|+|u_{2}|}=1,$ (4.32)
Here, in (4.29), we used that $|u_{1}|<|u^{\prime}|_{euc}$ which implies that $|u_{1}|>0$ or $|u^{\prime}|_{euc}-|u_{1}|>0$. In (4.30), (4.31), and (4.28), one can see that $u=u_{(2)}:=e_{2}$, $u=u_{(3)}:=e_{1}+e_{2}$, and $u=u_{(4)}:=e_{1}-e_{2}$ yield equality. We now define the functions $G_{1},G_{2},G_{3},G_{4}:\mathbb{R}^{d}\to[0,\infty]$ via
$G_{1}:=\frac{1}{\sqrt{2}}+\frac{\sqrt{2}-1}{\sqrt{2}}{\mathbbm{1}_{O\times(0,1)^{d-1}}},$
$G_{2}=\frac{1}{\sqrt{2}}+\frac{\sqrt{2}-1}{\sqrt{2}}\mathbbm{1}_{(0,1)^{d}},$
$G_{3}=G_{4}=1+(\sqrt{2}-1)\mathbbm{1}_{O\times(0,1)^{d-1}}.$ (4.33)
To summarise the above definitions more easily, note that for $i=1,\ldots,4$, we have that in $([0,1]^{d})^{c}\cup(O\times(0,1)^{d-1})$ it holds $G_{i}=|\nabla_{g}f_{i}|_{g}$ and in $((0,1)\setminus O)\times(0,1)^{d-1}$ we have that $G_{i}$ equals the respective upper bound established in (4.25)-(4.32).
Proposition 4.3.
For $1\leq i\leq 4$, it holds that $G_{i}$ is a weak upper gradient of $f_{i}$.
Proof.
Let $\gamma:[0,1]\to M$ be absolutely continuous and transversal to a $\mathcal{L}^{d}$-null set $N\supset N_{g}\cup N_{2}$. As $f_{i}\in C^{1}$, we may apply the chain rule. We have that for almost every $t\in[0,1]$, it holds that $\gamma_{t}\in([0,1]^{d})^{c}\cup(O\times(0,1)^{d-1})$ or $\gamma_{t}\in((0,1)\setminus O)\times(0,1)^{d-1}$. In the first case, we get that
$\Big{|}\frac{{\rm d}}{{\rm d}s}f_{i}\circ\gamma\Big{|}_{s=t}\Big{|}=|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|=|\langle\nabla_{g}f_{i},\dot{\gamma}_{t}\rangle_{g}|\leq|\nabla_{g}f_{i}|_{g}|\dot{\gamma}_{t}|_{g}=G_{i}(\gamma_{t})|\dot{\gamma}_{t}|.$ (4.34)
In the second case, we get that either $\dot{\gamma}_{t}=0$, in which case $\frac{{\rm d}}{{\rm d}s}f_{i}\circ\gamma\big{|}_{s=t}=0$ or
$\Big{|}\frac{{\rm d}}{{\rm d}s}f_{i}\circ\gamma\Big{|}_{s=t}\Big{|}=|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|=\frac{|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|}{\alpha(\dot{\gamma}_{t})}\alpha(\dot{\gamma}_{t})\leq G_{i}(\gamma_{t})|\dot{\gamma}_{t}|.$ (4.35)
Hence, for any transversal curve, (4.34) and (4.35) together yield that
$|f_{i}(\gamma_{1})-f_{i}(\gamma_{0})|\leq\int_{0}^{1}|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|\,{\rm d}t\leq\int_{0}^{1}G_{i}(\gamma_{t})|\dot{\gamma}_{t}|\,{\rm d}t,$ (4.36)
where in the last step we used (4.24) and (4.25)-(4.32). Now, we can conclude with Lemma 3.5. ∎
We next want to see that these are indeed the minimal weak upper gradients. For that we need the following result from [28, Lemma 4.23].
Lemma 4.4.
Let $k\in L^{1}_{\rm{loc}}(\mathbb{R}^{d},\mathcal{L}^{d})$, $x\in\mathbb{R}^{d}$, and $r>0$. For $\delta\in(0,r)\cap\mathbb{Q}$, $v\in B_{r}^{euc}(0)$, we define
$F_{v,\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}k(y+tv)\,{\rm d}\mathcal{L}^{n}(y),\quad\forall t,\;|t|<1.$
Then for $\mathcal{L}^{d}$-a.e. $x\in\mathbb{R}^{d}$ and all $\delta\in(0,r)\cap\mathbb{Q}$, $v\in B_{r}^{euc}(0)$, we have that $t=0$ is a Lebesgue point of $F_{v,\delta,x}$.
Proposition 4.5.
For $1\leq i\leq 4$, it holds that $G_{i}$ is the minimal weak upper gradient of $f_{i}$.
Proof.
We will argue similarly to the proof of Proposition 4.24 in [28]. Fix an $i\in\{1,2,3,4\}$. We argue by contradiction. Suppose that $G_{i}$ is not the minimal weak upper gradient of $f_{i}$. Then, there exists an $\varepsilon>0$ such that the set $D_{\varepsilon}:=\{x:|Df_{i}|_{w}(x)<G_{i}(x)-\varepsilon\}$ has positive $\mathcal{L}^{d}$-measure. Then we can choose a point $x\in D_{\varepsilon}\cap N_{2}^{c}$ such that $x$ is a Lebesgue point of $|Df_{i}|$, $G_{i}$ and such that $t=0$ is a Lebesgue point of
$F_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y),\ \mathrm{and}$
$G_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y),$
for all $\delta\in(0,1)\cap\mathbb{Q}$. This is possible by Lemma 4.4. Then we may choose $\delta\in(0,1)\cap\mathbb{Q}$ small enough such that
$\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}||Df_{i}|_{w}(y)-|Df_{i}|_{w}(x)|\,{\rm d}\mathcal{L}^{n}(y)\leq\frac{\varepsilon}{8},\ \mathrm{and}$ $\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|G_{i}(y)-G_{i}(x)|\,{\rm d}\mathcal{L}^{n}(y)\leq\frac{\varepsilon}{8}.$ (4.37)
Moreover, we can choose $\tau>0$ such that
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg{|}\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y)-\int_{B_{\delta}(x)}|Df_{i}|_{w}(y)\,{\rm d}\mathcal{L}^{n}(y)\Bigg{|}\,{\rm d}t\leq\frac{\varepsilon}{8},\ \mathrm{and}$
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg{|}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y)-\int_{B_{\delta}(x)}G_{i}(y)\,{\rm d}\mathcal{L}^{n}(y)\Bigg{|}\,{\rm d}t\leq\frac{\varepsilon}{8}.$ (4.38)
Then (4.37) and (4.2) together with the fact that $|Df_{i}|_{w}(x)<G_{i}(x)-\varepsilon$, yields that
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y){\rm d}t\geq\varepsilon-\frac{4\varepsilon}{8}=\frac{\varepsilon}{2}.$ (4.39)
We now define the curve $\gamma^{y,i}:[0,1]\to\mathbb{R}^{d}$ as $\gamma^{y,i}:t\mapsto y+2\tau(t-\frac{1}{2})u_{(i)}$. Then, we define the test plan
${\rm d}\boldsymbol{\pi}(\gamma):=\left\{\begin{array}[]{ll}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}{\rm d}\mathcal{L}^{d}(y)&\mathrm{if}\ \gamma=\gamma^{y,i},\ \mathrm{for\ some}\ y\in B_{\delta}(x),\\ 0&\mathrm{otherwise}.\end{array}\right.$ (4.42)
Now from the fact that $\mathsf{d}_{g}$ and $\mathsf{d}_{euc}$ are $\sqrt{2}$-equivalent, we get that for each $y\in\mathbb{R}^{d}$ and each $t\in[0,1]$ the metric speed $|\dot{\gamma}^{y,i}|\geq\frac{1}{\sqrt{2}}|u_{(i)}|_{euc}$.
$\int_{0}^{1}\int_{C^{0}([0,1],\mathbb{R}^{d})}(G_{i}(\gamma_{t})-|Df_{i}|_{w}(\gamma_{t}))|\dot{\gamma_{t}}|\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t$
$=\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}(G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)}))|\dot{\gamma}^{y,i}_{t}|\,{\rm d}\mathcal{L}^{n}(y){\rm d}t\geq\frac{\tau\varepsilon}{\sqrt{2}}.$ (4.43)
Now we observe that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(O\times(0,1)^{d-1})\cup([0,1]^{d})^{c}$, we have that $g\equiv Id$ a neighbourhood of $z$ and $\nabla_{g}f_{i}(z)=u_{(i)}$. Then by definition we have that $G_{i}(z)|\dot{\gamma}_{t}^{y,i}|=G_{i}|u_{(i)}|_{euc}=\langle\nabla_{euc}f_{i},u_{(i)}\rangle_{euc}=\partial_{u_{(i)}}f_{i}.$ Similarly, we have that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(N_{2}\cup N_{g})^{c}\setminus(O\times(0,1)^{d-1}\cup([0,1]^{d})^{c})$ that $\partial_{u_{(i)}}f_{i}(z)=G_{i}(z)|\dot{\gamma}_{t}^{y,i}|$ (see (4.35) and the equality cases of (4.25), (4.30), (4.31), and (4.32)).
Hence, we get that for every $y$ such that $\gamma^{y,i}$ is transversal to $N_{2}\cup N_{g}$, it holds
$f_{i}(\gamma^{y,i}_{1})-f_{i}(\gamma^{y,i}_{0})=\int_{0}^{1}\partial_{u_{(i)}}f_{i}(\gamma^{y,i}_{t})\,{\rm d}t=\int_{0}^{1}G_{i}(z)|\dot{\gamma}_{t}^{y,i}|\,{\rm d}t.$ (4.44)
On the other hand, Lemma 3.5 together with (4.2) and (4.44) yields that
$\int_{0}^{1}\int_{C^{0}([0,1],\mathbb{R}^{d})}|Df_{i}|_{w}(\gamma_{t})|\dot{\gamma_{t}}|\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t$
$\leq\int_{0}^{1}\int_{C^{0}([0,1],\mathbb{R}^{d})}G_{i}(\gamma_{t})|\dot{\gamma_{t}}|\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t-\frac{\tau\varepsilon}{\sqrt{2}}$
$=\int_{0}^{1}\int_{C^{0}([0,1],\mathbb{R}^{d})\cup\{\gamma\perp N_{2}\cup N_{g}\}}G_{i}(\gamma_{t})|\dot{\gamma_{t}}|\,{\rm d}\boldsymbol{\pi}(\gamma){\rm d}t-\frac{\tau\varepsilon}{\sqrt{2}}$
$<\int_{C^{0}([0,1],\mathbb{R}^{d})\cup\{\gamma\perp N_{2}\cup N_{g}\}}|f_{i}(\gamma_{1}^{y,i})-f_{i}(\gamma_{0}^{y,i})|\,{\rm d}\boldsymbol{\pi}(\gamma)=\int_{C^{0}([0,1],\mathbb{R}^{d})}|f_{i}(\gamma_{1}^{y,i})-f_{i}(\gamma_{0}^{y,i})|\,{\rm d}\boldsymbol{\pi}(\gamma).$
This contradicts the definition of a test plan and hence proves the Lemma. ∎
Now we have got all the ingredients to prove the following:
Proposition 4.6.
The metric measure space $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
Proof.
Pick a function $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi\equiv 1$ on $[-1,2]^{d}$. Consider the functions $\tilde{f}_{i}:=\phi f_{i}$ for $1\leq i\leq 4$. Then by the previous observations, we have that
$|D\tilde{f}_{i}|_{w}=G_{i}\ \mathrm{in}\ (-1,2)^{d}$
$|D\tilde{f}_{i}|_{w}=\frac{1}{\sqrt{2}}|\nabla_{euc}\tilde{f}_{i}|_{euc}\ \mathrm{in}\ ((-1,2)^{d})^{c}.$
More precisely, this gives
$|D\tilde{f}_{i}|_{w}=|\nabla_{g}\tilde{f}_{i}|_{g}\ \mathrm{a.e.\ in}\ (O\times(0,1)^{d-1})\cup((0,1)^{d})^{c},$ (4.45)
and
$|D\tilde{f}_{1}|_{w}=\frac{1}{\sqrt{2}},\ \mathrm{a.e.\ in}\ (0,1)^{d}\setminus O\times(0,1)^{d-1},$
$|D\tilde{f}_{2}|_{w}=|D\tilde{f}_{3}|_{w}=|D\tilde{f}_{4}|_{w}=1\ \mathrm{a.e.\ in}\ (0,1)^{d}\setminus(O\times(0,1)^{d-1}).$
Then it immediately follows that
$|D\tilde{f}_{3}|_{w}^{2}+|D\tilde{f}_{4}|_{w}^{2}=2(|D\tilde{f}_{1}|_{w}^{2}+|D\tilde{f}_{2}|_{w}^{2})\ \mathrm{a.e.\ in}\ (O\times(0,1)^{d-1})\cup((0,1)^{d})^{c},$ (4.46)
but
$|D\tilde{f}_{3}|_{w}^{2}+|D\tilde{f}_{4}|_{w}^{2}-2(|D\tilde{f}_{1}|_{w}^{2}+|D\tilde{f}_{2}|_{w}^{2})=-1\ \mathrm{a.e.\ in}\ (0,1)^{d}\setminus(O\times(0,1)^{d-1}).$ (4.47)
Moreover, we have that $\tilde{f}_{3}=\tilde{f}_{1}+\tilde{f}_{2}$ and $\tilde{f}_{4}=\tilde{f}_{1}-\tilde{f}_{2}$ and all four functions are in $W^{1,2}(\mathbb{R}^{d})$. But as $\mathcal{L}^{d}((0,1)^{d}\setminus O\times(0,1)^{d-1})\geq 1-2\kappa>0$ and the fact that $\mathrm{vol}_{g}$ and $\mathcal{L}^{d}$ are $2^{\frac{d}{2}}$-equivalent, we get that
$\mathrm{Ch}(\tilde{f}_{3})+\mathrm{Ch}(\tilde{f}_{4})\neq 2(\mathrm{Ch}(\tilde{f}_{1})+\mathrm{Ch}(\tilde{f}_{2})),$ (4.48)
which proves that the metric measure space $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian. ∎
4.3 An example of a Sobolev Riemannian metric, neither quasi-Riemannian nor infinitesimally Hilbertian
Fix a $d\geq 3$ and a $p\in(1,d-1)$. We will now refine the previous examples and construct a $d$-manifold $M$ and a metric $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}$, $g\in W^{1,p}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not quasi-Riemannian and such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
Define $\omega_{d-1}:=\mathcal{L}^{d-1}(B^{\mathbb{R}^{d-1}}_{1}(0))$. Pick a decreasing sequence $(R_{m})_{m\in\mathbb{N}}\subset(0,1)$ such that
$R_{m}\leq\frac{1}{2^{m+3}},$ (4.49)
$\sum_{m\in\mathbb{N}}2^{md}R_{m}^{d-1-p}<\infty,$ (4.50)
$\sum_{m\in\mathbb{N}}2^{(m+1)d+4}R_{m}<\frac{1}{4(1+\omega_{d-1})}.$ (4.51)
Fix a function $\eta\in C^{\infty}(\mathbb{R}^{d-1},[1,2])$ that satisfies $\eta|_{B_{1/2}(0)}\equiv 1$, $\eta|_{\mathbb{R}^{d-1}\setminus B_{3/4}(0)}\equiv 2$. We inductively define the finite sets $D_{m}\subset(0,1)^{d-1}$, open sets $O_{m}\subset(-1,2)^{d-1}$, a decreasing null sequence $(r_{m})_{m\in\mathbb{N}}\subset(0,1)$ and functions $\theta_{i}\in C^{\infty}(\mathbb{R}^{d-1})$ as follows:
$D_{0}:=O_{0}:=\emptyset,\ \theta_{0}\equiv 2\in C^{\infty}(\mathbb{R}^{d-1}),R_{0}:=\frac{1}{2},r_{0}:=\frac{1}{4}.$
Now for $m\geq 1$, define
$D_{m}:=\Big{(}\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}\cap(0,1)^{d-1}\Big{)}\setminus\overline{\big{(}\bigcup_{j\leq m-1}O_{j}\big{)}},$ (4.52)
$r_{m}:=\min\Big{(}R_{m},\frac{1}{2}\mathrm{dist}_{euc}(D_{m},\bigcup_{j\leq m-1}O_{j}),\frac{r_{m-1}}{2}\Big{)},$ (4.53)
$O_{m}:=\bigcup_{x\in D_{m}}B_{r_{m}}(x).$ (4.54)
Note that (4.53) together with (4.54) yield that
$\overline{O_{m}}\cap\overline{O_{j}}=\emptyset\ \mathrm{for\ all\ }j<m.$ (4.55)
Moreover, note that from $D_{m}\subset\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}$, together with (4.49) and (4.53), we have that for $x,x^{\prime}\in D_{m}$, $x\neq x^{\prime}$, it holds
$\overline{B_{r_{m}}(x)}\cap\overline{B_{r_{m}}(x^{\prime})}=\emptyset.$ (4.56)
Define the function $\psi_{m}\in C^{\infty}(\mathbb{R}^{d-1})$ via
$\psi_{m}(y):=\left\{\begin{array}[]{ll}2&\mathrm{if}\ y\in O_{m}^{c},\\ \eta\big{(}\frac{5(y-x)}{r_{m}}\big{)}&\mathrm{if}\ y\in B_{r_{m}}(x)\ \mathrm{for\ some\ }x\in D_{m}.\end{array}\right.$ (4.59)
By (4.54) and (4.56), we get that $\psi_{m}$ is well-defined. Now define
$\theta_{m}:=\min(\theta_{m-1},\psi_{m}).$ (4.60)
We observe that by (4.51), it holds
$\mathcal{L}^{d-1}\big{(}\bigcup_{m\in\mathbb{N}}\overline{O_{m}}\big{)}\leq\sum_{m=1}^{\infty}\#(\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}\cap(0,1)^{d-1})\mathcal{L}^{d-1}(B_{r_{m}}(0))\leq\omega_{d-1}\sum_{m=1}^{\infty}2^{(m+1)(d-1)+1}R_{m}^{d-1}\leq\frac{1}{4}.$ (4.61)
Define the set
$S:=(0,1)^{d-1}\setminus\bigcup_{m\in\mathbb{N}}\overline{O_{m}}.$ (4.62)
Lemma 4.7.
$\bigcup_{m\in\mathbb{N}}D_{m}$ is dense in $S$.
Proof.
Assume this was not the case. Then there exists a $y\in S$ and an $r>0$ such that $B_{r}(y)\cap\bigcup_{m\in\mathbb{N}}D_{m}=\emptyset$, where $B_{r}(y)=B^{\mathbb{R}^{d-1}}_{r}(y)$. By the monotonicity of the sequence $R_{m}$, there exists a $j\in\mathbb{N}$ such that
$R_{j^{\prime}}\leq\frac{r}{3},\ \mathrm{for}\ j^{\prime}\geq j.$ (4.63)
Moreover, as by definition $\bigcup_{k\leq j}\overline{O_{k}}\not\ni y$ is compact, there exists an $r^{\prime}\in(0,\frac{r}{3})$ such that
$B_{r^{\prime}}(y)\cap\bigcup_{k\leq j}\overline{O_{k}}=\emptyset.$ (4.64)
Now as $(\bigcup_{l\in\mathbb{N}}\frac{1}{2^{l+1}}\mathbb{Z}^{d-1}\setminus\bigcup_{l\leq j}\frac{1}{2^{l+1}}\mathbb{Z}^{d-1})\cap(0,1)^{d-1}$ is dense in $(0,1)^{d-1}$, can find $j<l\in\mathbb{N}$, $z\in\frac{1}{2^{l+1}}\mathbb{Z}^{d-1}\cap(0,1)^{d-1}$, such that $z\in B_{r^{\prime}}(y)$. Then by our assumption, we have that $z\notin\bigcup_{m\in\mathbb{N}}D_{m}$, hence $z\notin D_{l}$. Considering the $l$-th step of the above construction and (4.52), we get that $z\notin D_{l}$ implies
$z\in\bigcup_{k<l}\overline{O_{k}}.$
But then (4.64) implies that
$z\in\bigcup_{j<k<l}\overline{O_{k}}.$
Hence there exists a $j<k<l$ such that $z\in\overline{O_{k}}$, which together with (4.54) implies that $z\in B_{r_{k}}(x)$ for some $x\in D_{k}$. But then (4.63) and (4.64) give that
$|y-x|_{euc}\leq|z-y|_{euc}+|z-x|_{euc}\leq r_{k}+r^{\prime}\leq R_{j}+r^{\prime}<\frac{2r}{3}.$
But then $x\in B_{r}(y)\cap D_{k}$, which gives a contradiction. ∎
Now, we will prove the following useful property:
Lemma 4.8.
For each point $y\in\mathbb{R}^{d-1}\setminus\bigcup_{m\in\mathbb{N}}\overline{O_{m}}$, we have that
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r}=0\ \mathrm{and}$ (4.65)
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r^{d-1}}=0.$ (4.66)
Proof.
First, denote
$\zeta_{m}(y,r):=\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}.$ (4.67)
Fix an $m\in\mathbb{N}$. As $y\notin\overline{O_{m}}$, we get that for each $x\in D_{m}$, it holds $|x-y|_{euc}>r_{m}$, hence
$\mathrm{dist}_{euc}(y,\overline{B_{r_{m}/5}(x)})>\frac{4r_{m}}{5}.$ (4.68)
Thus, if $r\leq\frac{4r_{m}}{5}$, it follows that $\zeta_{m}(y,r)=0$. It now suffices to consider those $m$ such that $r_{m}\leq\frac{5}{4}r$. In that case, we get
$\zeta_{m}(y,r)$$\leq\#\big{\{}\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}\cap\overline{B_{r+\frac{r_{m}}{5}}(y)}\big{\}}\leq C(d)2^{(d-1)(m+1)}\frac{5^{d-1}}{4^{d-1}}r^{d-1},$
where $C(d)$ is a constant only depending on $d$. Then it follows that
[…]
Now the fact that $r_{m}>0$ together with (4.51) imply that the right-hand-side of (4.69) converges to 00 as $r\to 0$, which implies (4.66). Moreover,
[…]
which using (4.51) once again yields (4.65). ∎
Proposition 4.9.
The functions $\theta_{m}$ converge to a function $\theta$ in $W^{1,p}_{\rm loc}(\mathbb{R}^{d-1})$ as $m\to\infty$.
Proof.
We first note that for all $m$, it holds $\theta_{m}\equiv 2$ outside $(-1,2)^{d-1}$, hence, it suffices to consider its behaviour inside $(-1,2)^{d-1}$. Note that (4.59) implies that $\psi_{m}^{-1}([1,2))\subset O_{m}$. Then (4.55) together with (4.60) yields that for any $m^{\prime}<m$ it holds
$\theta_{m}=\theta_{m^{\prime}}+\sum_{j=m^{\prime}+1}^{m}(\psi_{j}-2).$ (4.71)
Then, as $(\psi_{j}-2)\in C_{c}^{\infty}((-1,2)^{d-1})$, the Poincaré inequality together with (4.56) yields that
$\norm{(\psi_{j}-2)}^{p}_{W^{1,p}(\mathbb{R}^{d-1})}$ $\leq C\int_{(-1,2)^{d-1}}|\nabla_{euc}\psi_{j}|^{p}\,{\rm d}\mathcal{L}^{d-1}$
$=C|D_{j}|\int_{B_{r_{j}}(0)}\frac{1}{r_{j}^{p}}|\nabla_{euc}\eta|^{p}\Big{(}\frac{x}{r_{j}}\Big{)}\,{\rm d}\mathcal{L}^{d-1}(x)$
$\leq C2^{(j+1)(d-1)}r^{-p}_{j}\mathcal{L}^{d-1}(B_{r_{j}}(0))|\nabla_{euc}\eta|^{p}_{L^{\infty}}$
$\leq C2^{(j+1)(d-1)}r^{d-1-p}_{j}.$ (4.72)
Hence, for $m^{\prime}\leq m$, we have that
$\norm{\nabla_{euc}\theta_{m}-\nabla_{euc}\theta_{m^{\prime}}}_{L^{p}((-1,2)^{d-1})}^{p}\leq C\sum_{j=m^{\prime}+1}^{m}2^{(j+1)(d-1)}r^{d-1-p}_{j}\leq C\sum_{j=m^{\prime}+1}^{\infty}2^{(j+1)(d-1)}r^{d-1-p}_{j}.$
Now, (4.50) yields that the right hand side goes to 00 as $m^{\prime}\to\infty$, thus, $(\nabla_{euc}\theta_{m})_{m}$ is a null-sequence in $L^{p}(\mathbb{R}^{d})$. As for all $m$ one has that $\theta_{m}-2\in C_{c}^{\infty}((-1,2)^{d-1},\mathbb{R})$, the Poincaré inequality yields that $\theta_{m}$ converges to a function $\theta$ in $W^{1,p}_{\rm loc}(\mathbb{R}^{d-1})$. ∎
Our inductive construction also implies that $\theta_{m}$ converges pointwise and is bounded in $L^{\infty}$, hence, we can describe $\theta$ explicitly: For all $y\in\mathbb{R}^{d-1}\setminus\bigcup_{m\in\mathbb{N}}\overline{O_{m}}$, we have that $\theta_{m}(y)=2$ for all $m\in\mathbb{N}$. For all $y\in\bigcup_{m\in\mathbb{N}}\overline{O_{m}}$, there exists a $m^{*}$ such that $y\in\overline{O_{m^{*}}}$. Then for any $m\geq m^{*}$, (4.55) guarantees that $\theta_{m}(y)=\theta_{m^{*}}(y)$. Now as by the dominated convergence theorem, $\theta$ is the pointwise limit of the $\theta_{m}$, we can conclude that
$1\leq\theta\leq 2\ \mathrm{on}\ \mathbb{R}^{d-1},$ (4.73)
$\theta\equiv 2\ \mathrm{on}\ \mathbb{R}^{d-1}\setminus\bigcup_{m\in\mathbb{N}}\overline{O_{m}},$ (4.74)
$\theta^{-1}([1,2))\subset\bigcup_{m\in\mathbb{N}}\bigcup_{x\in D_{m}}B_{r_{m}/5}(x).$ (4.75)
Moreover, we get that for each $m\in\mathbb{N}$, $x\in D_{m}$, it holds
$\theta\equiv 1\ \mathrm{on}\ B_{r_{m}/10}(x).$ (4.76)
Proposition 4.10.
The set $N_{\theta}\subset\mathbb{R}^{d-1}$ of non-Lebesgue points of $\theta$ can be chosen as a subset of the $\mathcal{L}^{d-1}$-null set $\bigcup_{m\in\mathbb{N}}\partial O_{m}$.
Proof.
For each point $y\in\mathbb{R}^{d-1}\setminus\bigcup_{m\in\mathbb{N}}\overline{O_{m}}$, we can apply (4.66) and get that
$\frac{1}{\mathcal{L}^{d-1}(B_{r}(y))}\int_{B_{r}(y)}|\theta(x)-2|{\rm d}\mathcal{L}^{d-1}$ $\leq\frac{\mathcal{L}^{d-1}(B_{r}(y)\cap\bigcup_{m\in\mathbb{N}}\bigcup_{x\in D_{m}}B_{r_{m}/5}(x))}{\mathcal{L}^{d-1}(B_{r}(y))}$
$\leq C\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r^{d-1}}\to 0\ \mathrm{as}\ r\to 0.$
For each point $y\in\bigcup_{m\in\mathbb{N}}O_{m}$, we have that $\theta$ is continuous in a neighbourhood of $y$, hence $y$ is a Lebesgue point. For each $m$, we have that $\partial O_{m}$ is a $\mathcal{L}^{d-1}$-null set, hence so is their countable union. ∎
We are now ready to define the metric $g$ on $M:=\mathbb{R}^{d}$ via
$g(x^{\prime},x_{d})=\theta(x^{\prime})Id,\ x^{\prime}\in\mathbb{R}^{d-1},x_{d}\in\mathbb{R}.$ (4.77)
Again $(M,g)$ satisfies (ii) from Assumption 3.1 and $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1).
4.3.1 Analysis of quasi-Riemannianity
We now have enough information to prove Theorem 1. Note that by (4.61) and (4.62), we have that $S\subset(0,1)^{d-1}$ satisfies $\mathcal{L}^{d-1}(S)>0$ and $\theta=2$ on $S$ (see Proposition 4.10). Then consider $E:=S\times(0,\frac{1}{2})\subset M$. Then $\mathrm{vol}_{g}(E)\geq\mathcal{L}^{d}(E)>0$. For $y\in E$, define the curve $\gamma^{y}:[0,1]\to(0,1)^{d-1}\times(-2,2),t\mapsto y+te_{d}$ and define the test plan $\boldsymbol{\pi}\in\mathcal{P}(C([0,1],\mathbb{R}^{d}))$ via
${\rm d}\boldsymbol{\pi}(\gamma):=\left\{\begin{array}[]{ll}\frac{1}{\mathcal{L}^{d}(E)}{\rm d}\mathcal{L}^{d}(y)&\mathrm{if}\ \gamma=\gamma^{y},\ \mathrm{for\ some}\ y\in E,\\ 0&\mathrm{otherwise}.\end{array}\right.$
Now the strategies from Subsection 4.1 together with Lemma 4.7 and Proposition 4.10 imply that for all $y\in E$, and almost every $t\in[0,1]$, it holds $|\dot{\gamma}^{y}_{t}|=1<\sqrt{2}=|\dot{\gamma}^{y}_{t}|_{g}$. Hence, $\boldsymbol{\pi}$ is supported on curves for which the metric speed is strictly smaller than the norm of the derivative, for almost every time.
Now we can consider the function $f\in C^{1}(M)$ given by $f:(x^{\prime},x_{d})\to x_{d}$. Then $f$ has slope $|Df|=1$ on $S\times\mathbb{R}$, but $|\nabla_{g}f|_{g}<1$ on $S\times\mathbb{R}$. Finally, $f$ and $\boldsymbol{\pi}$ together yield that $|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive measure, which proves Theorem 1.
We will conclude this observation with the following remark.
Remark 4.11.
From the Sobolev embedding theorem, we get that for a $d$-dimensional manifold with a Riemannian metric $g\in W^{1,p}_{\rm loc}$ for some $p>d$, that $g$ is automatically continuous. Then the results we recalled in Subsection 2.2, yield that $(M,\mathsf{d}_{g},g)$ satisfies condition (iii) from Definition 3.6 for $C^{1}$-functions. Theorem 1 shows that this is no longer given if $p<d-1$ and therewith works towards solving Problem 2 in [10], Chapter 3: spazi metrici quasi-riemanniani: the question what regularity assumption on $g$ is necessary such that the $\mathsf{d}_{g}$-slope of a Lipschitz function equals the $g$-norm of the gradient almost everywhere (see also (2.2), Theorem (2.18), and Theorem (6.1) in [9]).
4.3.2 Analysis of infinitesimal Hilbertianity
From here, we will analyse the infinitesimal structure of the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ and prove Theorem 2. Pick a point $y=(y^{\prime},y_{d})\in M$ such that $y^{\prime}\in S$. We will work in a neighbourhood around $y$ and restrict to local coordinates. Pick a vector $u\in\mathbb{R}^{d}$. We want to understand the metric speed at time 00 of the curve $\gamma^{y,u}:[0,\xi]\to\mathbb{R}^{d},t\mapsto y+tu$, $\xi>0$. Define the null set
$N_{g}:=(\partial[0,1]^{d-1}\cup\bigcup_{m\in\mathbb{N}}\partial O_{m})\times\mathbb{R}\subset M.$ (4.78)
Fix an $\varepsilon>0$ and find an $r_{\varepsilon}>0$ such that for $r\in(0,r_{\varepsilon})$, the expressions in (4.65) and (4.66) are smaller than or equal to $\varepsilon$. Now for some $0<r<r_{\varepsilon}$ choose $z\in M$ such that $\mathsf{d}_{euc}(z,y)=\frac{r}{12}$, which implies that any $\frac{4}{3}$-shortest curve transversal to $N_{g}$ (i.e. a curve $\gamma\perp N_{g}$ from $y$ to $z$, such that $L_{g}(\gamma)\leq\frac{4}{3}\mathsf{d}_{g}(y,z)$) from $y$ to $z$ lies inside $B_{r}^{euc}(y)$. Pick such a curve $\gamma$ and assume it to be $1$-$\mathsf{d}_{euc}$-Lipschitz (which can be done if $r$ is small enough, as $\mathsf{d}_{g}$ and $\mathsf{d}_{euc}$ are $\sqrt{2}$-equivalent).
Denote
$\Omega_{m}:=\bigcup_{x\in D_{m}}B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)\subset(-1,2)^{d-1}.$ (4.79)
We first note that by the definition of the metric $g$ it holds that
$L_{g}(\gamma)\geq\int_{\gamma^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t+\int_{\gamma^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}_{t}|_{euc}\,{\rm d}t.$ (4.80)
Now observe that
$\gamma^{-1}\Big{(}\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}}\times\mathbb{R}\Big{)}=\bigcup_{m\in\mathbb{N}}\gamma^{-1}(\overline{\Omega_{m}}\times\mathbb{R}).$
Hence, we can pick a $J\in\mathbb{N}$ such that
$\mathcal{L}^{1}\Big{(}\gamma^{-1}\Big{(}\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}}\times\mathbb{R}\Big{)}\Big{)}\leq\frac{r\varepsilon}{3}+\mathcal{L}^{1}\Big{(}\gamma^{-1}\Big{(}\bigcup_{m\leq J}\overline{\Omega_{m}}\times\mathbb{R}\Big{)}\Big{)}.$ (4.81)
For any $x\in\bigcup_{m\in\mathbb{N}}D_{m}$, write $m(x)$ as the natural number such that $x\in D_{m(x)}$. Now the set $\bigcup_{m\leq J}D_{m}$ is finite so we may write it as $\{x_{l}\}_{l=1}^{k}\subset\mathbb{R}^{d-1}$ for some $k\in\mathbb{N}$. Write $m_{l}:=m(x_{l})$. We will inductively define Lipschitz-continuous curves $\gamma^{(l)}:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ for $l=0,\ldots,k$ as follows: Define $\gamma^{(0)}:=\gamma$. Now for $l\geq 1$ define
$t_{l}^{\min}:=\min(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{euc}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}),$
$t_{l}^{\max}:=\max(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{euc}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}).$
Define
$w_{l}:=\int_{t_{l}^{\min}}^{t_{l}^{\max}}\dot{\gamma}^{(l-1)}_{t}\,{\rm d}t\in\mathbb{R}^{d},$ (4.82)
and
$\gamma^{(l)}_{t}:=\left\{\begin{array}[]{ll}\gamma^{(l-1)}_{t}&\mathrm{if}\ t\in[0,t_{l}^{\min}],\\ \gamma^{(l-1)}_{t_{l}^{\min}}+\frac{t-t_{l}^{\min}}{t_{l}^{\max}-t_{l}^{\min}}w_{l}&\mathrm{if}\ t\in[t_{l}^{\min},t_{l}^{\max}],\\ \gamma^{(l-1)}_{t}&\mathrm{if}\ t\in[t_{l}^{\max},1].\end{array}\right.$
Claim 1. For each $l=1,\ldots,k$, $\gamma^{(l)}$ satisfies
$\int_{(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t+\int_{(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t$
$\geq\int_{(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(l)}_{t}|_{euc}\,{\rm d}t+\int_{(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l)}_{t}|_{euc}\,{\rm d}t.$ (4.83)
Proof of claim 1.
This can be seen by noticing that
$(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})=(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})\setminus[t_{l}^{\min},t_{l}^{\max}]\ \mathrm{and}$
$(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})=(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})\cup[t_{l}^{\min},t_{l}^{\max}].$ (4.84)
Moreover, $\gamma^{(l-1)}=\gamma^{(l)}$ on $[t_{l}^{\min},t_{l}^{\max}]^{c}$, thus
$\int_{[t_{l}^{\min},t_{l}^{\max}]^{c}\cap(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t+\int_{[t_{l}^{\min},t_{l}^{\max}]^{c}\cap(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t$
$=\int_{[t_{l}^{\min},t_{l}^{\max}]^{c}\cap(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(l)}_{t}|_{euc}\,{\rm d}t+\int_{[t_{l}^{\min},t_{l}^{\max}]^{c}\cap(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l)}_{t}|_{euc}\,{\rm d}t,$ (4.85)
and
$\int_{[t_{l}^{\min},t_{l}^{\max}]\cap(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t+\int_{[t_{l}^{\min},t_{l}^{\max}]\cap(\gamma^{(l-1)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l-1)}_{t}|_{euc}\,{\rm d}t$
$\geq\int_{[t_{l}^{\min},t_{l}^{\max}]}|\dot{\gamma}_{t}^{(l-1)}|_{euc}\,{\rm d}t\geq|w_{l}|_{euc}=\int_{[t_{l}^{\min},t_{l}^{\max}]\cap(\gamma^{(l)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(l)}_{t}|_{euc}\,{\rm d}t.$ (4.86)
The sum of (4.3.2) and (4.3.2) yields the claim. ∎
Furthermore, we observe that (4.3.2) yields
$\mathcal{L}^{1}((\gamma^{(l)})^{-1}((\bigcup_{m>J}\overline{\Omega_{m}})\times\mathbb{R}))$ $=\mathcal{L}^{1}((\gamma^{(l-1)})^{-1}((\bigcup_{m>J}\overline{\Omega_{m}})\times\mathbb{R})\setminus[t_{l}^{\min},t_{l}^{\max}])$
$\leq\mathcal{L}^{1}((\gamma^{(l-1)})^{-1}((\bigcup_{m>J}\overline{\Omega_{m}})\times\mathbb{R})),$ (4.87)
for all $l=1,\ldots,k$.
Claim 2. For each $l=1,\ldots,k$, $l^{\prime}\leq l$, $(\gamma^{(l)})^{-1}(\overline{B_{r_{m_{l^{\prime}}}/5}(x_{l^{\prime}})}\times\mathbb{R})$ is a closed interval.
Proof of claim 2..
We will prove this by induction. For $l=1$ the claim is true as $(\gamma^{(1)})^{-1}(\overline{B_{r_{m_{1}}/5}(x_{1})}\times\mathbb{R})=[t_{1}^{\min},t_{1}^{\max}]$. Now assume the claim holds true for some $l<k$. Now we have that $(\gamma^{(l+1)})^{-1}(\overline{B_{r_{m_{l+1}}/5}(x_{l+1})}\times\mathbb{R})=[t_{l+1}^{\min},t_{l+1}^{\max}]$. For $l^{\prime}\leq l$ denote by $I^{(l)}_{l^{\prime}}:=(\gamma^{(l)})^{-1}(\overline{B_{r_{m_{l^{\prime}}}/5}(x_{l^{\prime}})}\times\mathbb{R})$. By the construction of $\gamma^{(l+1)}$, (4.55), and (4.56), we have that $t_{l+1}^{\min},t_{l+1}^{\max}\notin I^{(l)}_{l^{\prime}}$. Hence, as $I^{(l)}_{l^{\prime}}$ is a closed interval by the induction hypothesis, it holds that $[t_{l+1}^{\min},t_{l+1}^{\max}]\supset I^{(l)}_{l^{\prime}}$, or $[t_{l+1}^{\min},t_{l+1}^{\max}]\cap I^{(l)}_{l^{\prime}}=\emptyset$. In the first case, we get that $I^{(l+1)}_{l^{\prime}}=\emptyset$ and in the second one it holds $I^{(l+1)}_{l^{\prime}}=I^{(l)}_{l^{\prime}}$. Both are closed intervals, which finishes the induction step and proves the claim. ∎
Finally, we note that for all $l=0,\ldots,k$, $\gamma^{(l)}$ is $1$-$\mathsf{d}_{euc}$-Lipschitz. This is true by assumption for $l=0$. For $l>0$ this follows inductively, using that in $\mathbb{R}^{d}$, $w_{l}=\gamma^{(l-1)}_{t^{\max}_{l}}-\gamma^{(l-1)}_{t^{\min}_{l}}$, hence by the induction hypothesis, $|w_{l}|_{euc}\leq t^{\max}_{l}-t^{\min}_{l}$.
Now we can combine (4.80) with (4.3.2) to get that
$L_{g}(\gamma)\geq\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(k)}_{t}|_{euc}\,{\rm d}t+\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(k)}_{t}|_{euc}\,{\rm d}t.$
Now (4.81) and (4.3.2), together with the fact that $\gamma^{(k)}$ is $1$-$\mathsf{d}_{euc}$-Lipschitz continuous yield that
$L_{g}(\gamma)+r\varepsilon\geq\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\leq J}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\sqrt{2}|\dot{\gamma}^{(k)}_{t}|_{euc}\,{\rm d}t+\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\leq J}\overline{\Omega_{m}})\times\mathbb{R})}|\dot{\gamma}^{(k)}_{t}|_{euc}\,{\rm d}t.$ (4.88)
As in the previous example, we define
$v_{(1)}:=\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\leq J}\overline{\Omega_{m}})\times\mathbb{R})}\dot{\gamma}^{(k)}\,{\rm d}t\in\mathbb{R}^{d},$
$v_{(2)}:=\int_{(\gamma^{(k)})^{-1}((\bigcup_{m\leq J}\overline{\Omega_{m}})^{c}\times\mathbb{R})}\dot{\gamma}^{(k)}\,{\rm d}t\in\mathbb{R}^{d}.$
Then,
$v_{(1)}+v_{(2)}=z-y.$
Applying classical facts on the Euclidean space to (4.88) yields that
$L_{g}(\gamma)+r\varepsilon\geq\sqrt{2}|v_{(2)}|_{euc}+|v_{(1)}|_{euc}.$ (4.89)
Let $P^{(d-1)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the orthogonal projection mapping $\mathbb{R}^{d-1}\times\mathbb{R}\ni(x^{\prime},x_{d})\mapsto(x^{\prime},0)$. Now by Claim 2 together with our choice of $r_{\varepsilon}$ (see (4.65)), we get that
$|P^{(d-1)}v_{(1)}|_{euc}$ $=\Big{|}\sum_{l=1}^{k}\int_{(\gamma^{(k)})^{-1}(\overline{B_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big{|}_{euc}$
$\leq\sum_{l=1}^{k}\Big{|}\int_{(\gamma^{(k)})^{-1}(\overline{B_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big{|}_{euc}$
$\leq\sum_{l=1}^{k}\mathrm{diam}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})})$
$\leq 2r\frac{\sum_{m\in\mathbb{N}}\frac{r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{r}(y^{\prime})}\}|}{r}\leq 2r\varepsilon,$ (4.90)
where we wrote $y=(y^{\prime},y_{d})\in\mathbb{R}^{d-1}\times\mathbb{R}$. Now define
$w_{(1)}:=v_{(1)}-P^{(d-1)}(v_{(1)})=(v_{(1)})_{d}e_{d}\in\mathrm{span}(e_{d})\subset\mathbb{R}^{d},\ w_{(2)}=v_{(2)}+P^{(d-1)}(v_{(1)})\in\mathbb{R}^{d}.$
Then $w_{(1)}+w_{(2)}=z-y$ and
$|w_{(1)}-v_{(1)}|_{euc},|w_{(2)}-v_{(2)}|_{euc}\leq 2r\varepsilon.$
Hence, with (4.89), we get that
$L_{g}(\gamma)+6r\varepsilon\geq\sqrt{2}|w_{(2)}|_{euc}+|w_{(1)}|_{euc},$ (4.91)
for vectors $w_{(1)},w_{(2)}$ with $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{d})$. We want to minimise the right hand side of (4.91) under the thereafter mentioned constraints. This will be done similarly as in the previous example (see from (4.14) onwards).
Write again $z-y=u=(u^{\prime},u_{d})\in\mathbb{R}^{d-1}\times\mathbb{R}$. By potentially applying a linear coordinate transform from $SO(d-1)$, we may assume that $u^{\prime}\in\mathrm{span}(e_{1})$. Then we reduce the right-hand-side of (4.91) by orthogonally projecting $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{d})$. Now if $u\in\mathrm{span}(e_{1})$, we get that $w_{(1)}=0$ and hence
$L_{g}(\gamma)+6r\varepsilon\geq\sqrt{2}|z-y|_{euc}.$
Otherwise, we have that $u_{d}\neq 0$. Then $u$ can be written as $a(be_{1}+e_{d})$. Then writing $w_{(1)}=a(1-c)e_{d}$ and $w_{(2)}=a(be_{1}+ce_{d})$ reduces to minimising $f_{b}(c)$ as defined in (4.16). Together with $r=12|u|_{euc}$, this yields
$L_{g}(\gamma)\geq\left\{\begin{array}[]{ll}|u_{d}|+|u^{\prime}|_{euc}-100\varepsilon|u|_{euc}&\mathrm{if}\ |u_{d}|\geq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-100\varepsilon|u|_{euc}&\mathrm{otherwise}.\end{array}\right.$
As before, together with $\mathsf{d}_{g}\geq\boldsymbol{\delta}_{N_{g}}$ this yields
$\mathsf{d}_{g}(y,y+u)\geq\left\{\begin{array}[]{ll}|u_{d}|+|u^{\prime}|_{euc}-100\sigma(y,u)|u|_{euc}&\mathrm{if}\ |u_{d}|\geq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}-100\sigma(y,u)|u|_{euc}&\mathrm{otherwise},\end{array}\right.$ (4.94)
where
$\sigma(y,u):=\frac{\sum_{m\in\mathbb{N}}\frac{2r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{|u|_{euc}}(y^{\prime})}\}|}{|u|_{euc}}\to 0,\ \mathrm{as}\ |u|_{euc}\to 0,$
as of Lemma 4.8.
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ and work with the regularisation. Fix $\epsilon>0$. For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2Id$. Hence,
$\mathsf{d}_{g_{\varsigma}}(y,z)\leq\sqrt{2}|u|_{euc}.$
By Lemma 4.7, we can find an $m\in\mathbb{N}$ and an $x\in D_{m}$ such that $|P^{(d-1)}(y)-(x,0)|\leq\epsilon|u|_{euc}$. Now for $\varsigma\in(0,r_{m}/20)$, it holds $g_{\varsigma}=1$ on $\{x\}\times\mathbb{R}$. If $|u_{d}|>|u^{\prime}|_{euc}$, write $b=\frac{|u^{\prime}|_{euc}}{|u_{d}|}$ and note that the concatenation of the curves
$\gamma_{1}:[0,1]\to\mathbb{R}^{d},t\mapsto y+t((x,0)-P^{(d-1)}y),$
$\gamma_{2}:[0,1]\to\mathbb{R}^{d},t\mapsto y+((x,0)-P^{(d-1)}y)+t(1-b)u_{d}e_{d},$
$\gamma_{3}:[0,1]\to\mathbb{R}^{d},t\mapsto y+((x,0)-P^{(d-1)}y)+(1-b)u_{d}e_{d}-t((x,0)-P^{(d-1)}y),\ \mathrm{and}$
$\gamma_{4}:[0,1]\to\mathbb{R}^{d},t\mapsto y+(1-b)u_{d}e_{d}+t((u^{\prime},0)+bu_{d}e_{d}),$
connects $y$ with $z$ and has $g_{\varsigma}$ length at most $|u^{\prime}|_{euc}+|u_{d}|_{euc}+3\epsilon$. As $\epsilon$ was arbitrary, we get that
$\mathsf{d}_{g}(y,y+u)=\lim_{\varsigma\to 0}\mathsf{d}_{g_{\varsigma}}(y,y+u)\leq\alpha(u):=\left\{\begin{array}[]{ll}|u_{d}|+|u^{\prime}|_{euc}&\mathrm{if}\ |u_{d}|\geq|u^{\prime}|_{euc},\\ \sqrt{2}|u|_{euc}&\mathrm{otherwise}.\end{array}\right.$ (4.97)
Together with Lemma 3.3, this gives that for any $y\in S\times\mathbb{R}$, the metric speed of a curve $\gamma:[0,1]\to M$, $t\in[0,1]$ such that $\dot{\gamma}_{t}=:u$ exists and $\gamma_{t}=y$ is given by $\alpha(u)$.
Now the argument works similarly as in the previous example: We note that for any $y\in(([0,1]^{d-1})^{c}\times\mathbb{R})\cup(\bigcup_{m\in\mathbb{N}}O_{m}\times\mathbb{R})$ we have that $g$ is continuous in a neighbourhood of $y$ and the metric speed of a curve is given by the $g$-norm of its derivative, wherever it exists. By the results from [28], the minimal weak upper gradient of any $C^{1}$-function $f$ is then also almost everywhere given by the $g$-norm of its $g$-gradient $\nabla_{g}f$. Let $\tilde{\phi}\in C_{c}^{\infty}(\mathbb{R}^{d-1})$ such that $\tilde{\phi}((-2,2)^{d-1})=\{1\}$ and $\tilde{\phi}\equiv 0$ on $((-3,3)^{d-1})^{c}$. Define $\phi\in C^{\infty}(M)$ locally via $\phi:\mathbb{R}^{d-1}\times\mathbb{R},(x^{\prime},x_{d})\mapsto\tilde{\phi}(x^{\prime})$. Now consider the functions
$f_{1}:x\mapsto\phi(x)x_{1},f_{2}:x\mapsto\phi(x)|x_{d}|,f_{3}=f_{1}+f_{2},f_{4}=f_{1}-f_{2}.$
Define the functions $G_{i}$ for $i=1,\ldots,4$ $\mathrm{vol}_{g}$-almost everywhere (more precisely on $M\setminus N_{g}$) via
$G_{1}(z):=\left\{\begin{array}[]{ll}\frac{1}{\sqrt{2}}&\mathrm{if}\ z\in S\times\mathbb{R},\\ |\nabla_{g}f_{1}|_{g}&\mathrm{otherwise}.\end{array}\right.,G_{2}(z):=\left\{\begin{array}[]{ll}1&\mathrm{if}\ z\in S\times\mathbb{R},\\ |\nabla_{g}f_{2}|_{g}&\mathrm{otherwise}.\end{array}\right.$
$G_{3}(z):=\left\{\begin{array}[]{ll}1&\mathrm{if}\ z\in S\times\mathbb{R},\\ |\nabla_{g}f_{3}|_{g}&\mathrm{otherwise}.\end{array}\right.,G_{4}(z):=\left\{\begin{array}[]{ll}1&\mathrm{if}\ z\in S\times\mathbb{R},\\ |\nabla_{g}f_{4}|_{g}&\mathrm{otherwise}.\end{array}\right.$
As in Proposition 4.5, one can show that $G_{i}$ is the minimal weak upper gradient of $f_{i}$ for $i=1,\ldots,4$. Now from the above definition, one can see that $f_{i},G_{i}$ are in $L^{\infty}_{\rm loc}(M)$. From the definition, we have that outside $S\times\mathbb{R}$, we have that $2G_{1}^{2}+2G_{2}^{2}=G_{3}^{2}+G_{4}^{2}$ almost everywhere. Inside $S\times\mathbb{R}$, we have that
$2G_{1}^{2}+2G_{2}^{2}-(G_{3}^{2}+G_{4}^{2})=1\neq 0$ (4.98)
As in the previous example, this is the main fact that will yield non-infinitesimal Hilberianity. We only have to navigate around the problem that $G_{i}\notin L^{2}(\mathbb{R}^{d},\mathrm{vol}_{g})$. In a final step define a function $\eta\in C_{c}^{\infty}(\mathbb{R})$ such that $\eta(x)=\eta(-x)$ for all $x\in\mathbb{R}$ and such that $\eta\equiv 1$ on $(-1,1)$ and $\eta\equiv 0$ on $(-2,2)^{c}$. For $n\geq 1$ define $\eta_{n}\in C_{c}^{\infty}(\mathbb{R})$ via $\eta_{n}\equiv 1$ on $(-n,n)$, $\eta_{n}(x)=\eta(|x|-n+1)$ for $|x|\geq n$ and for $1\leq i\leq 4$, define the functions $f_{i,n}:\mathbb{R}^{d}\to\mathbb{R}$ via
$f_{1,n}:(x_{1},\ldots,x_{d})=x\mapsto\eta_{n}(x_{d})\phi(x)x_{1},$
$f_{2,n}:(x_{1},\ldots,x_{d})=x\mapsto\eta_{n}(x_{d})\phi(x)(|x_{d}|-n+1),$
$f_{3,n}:=f_{1,n}+f_{2,n},\ f_{4,n}:=f_{1,n}-f_{2,n}.$
Then $\mathrm{supp}\,f_{i,n}\subset[-3,3]^{d-1}\times[-n-1,n+1]$. Now, with similar methods as before, using the fact that the minimal weak upper gradient depends locally on the function, we get that the minimal weak upper gradient $|Df_{i,n}|_{w}$ of $f_{i,n}$ can be described via
$|Df_{i,n}|_{w}(z):=\left\{\begin{array}[]{ll}G_{i}(z)&\mathrm{if}\ z\in S\times(-n,n),\\ |\nabla_{g}f_{i,n}|_{g}(z)&\mathrm{if}\ z\in((-3,3)^{d-1}\setminus S)\times(-n,n),\\ |Df_{i,1}|_{w}(z-(n-1)e_{d})&\mathrm{if}\ z\in(-3,3)^{d-1}\times[n,n+1],\\ |Df_{i,1}|_{w}(z+(n-1)e_{d})&\mathrm{if}\ z\in(-3,3)^{d-1}\times[-n-1,-n],\\ 0&\mathrm{otherwise}.\end{array}\right.$
Now,
$\int_{M}2|Df_{1,n}|^{2}_{w}+2|Df_{2,n}|^{2}_{w}-|Df_{3,n}|^{2}_{w}-|Df_{4,n}|^{2}_{w}\,{\rm d}\mathrm{vol}_{g}$
$=\int_{(-3,3)^{d-1}\times(-n,n)}2|Df_{1,n}|^{2}_{w}+2|Df_{2,n}|^{2}_{w}-|Df_{3,n}|^{2}_{w}-|Df_{4,n}|^{2}_{w}\,{\rm d}\mathrm{vol}_{g}$
$+\int_{(-3,3)^{d-1}\times([-n-1,-n]\cup[n,n+1])}2|Df_{1,n}|^{2}_{w}+2|Df_{2,n}|^{2}_{w}-|Df_{3,n}|^{2}_{w}-|Df_{4,n}|^{2}_{w}\,{\rm d}\mathrm{vol}_{g}$
$=n\int_{S\times(-1,1)}2G_{1}^{2}+2G_{2}^{2}-G_{3}^{2}-G_{4}^{2}\,{\rm d}\mathrm{vol}_{g}$
$+\int_{(-3,3)^{d-1}\times((-2,-1)\cup(1,2))}2|Df_{1,1}|^{2}_{w}+2|Df_{2,1}|^{2}_{w}-|Df_{3,1}|^{2}_{w}-|Df_{4,1}|^{2}_{w}\,{\rm d}\mathrm{vol}_{g}$
$=2n\mathrm{vol}_{g}(S\times(0,1))+\int_{(-3,3)^{d-1}\times((-2,-1)\cup(1,2))}2|Df_{1,1}|^{2}_{w}+2|Df_{2,1}|^{2}_{w}-|Df_{3,1}|^{2}_{w}-|Df_{4,1}|^{2}_{w}\,{\rm d}\mathrm{vol}_{g},$
where in the last two steps, we first used the fact that $02|\nabla_{g}f_{1,n}|_{g}^{2}+2|\nabla_{g}f_{2,n}|_{g}^{2}-|\nabla_{g}f_{3,n}|_{g}^{2}-|\nabla_{g}f_{4,n}|_{g}^{2}=$ 0 for all $n\in\mathbb{N}$ and then applied (4.98). We have that $|Df_{i,1}|^{2}_{w}\in L^{\infty}$ and hence the last integral in the above equation is finite, so as $\mathrm{vol}_{g}(S\times(0,1))>0$, we get that
$\lim_{n\to\infty}2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})\infty.$ (4.99)
In particular, there exists an $n\in\mathbb{N}$ such that $2\mathrm{Ch}(f_{1,n})+2\mathrm{Ch}(f_{2,n})-\mathrm{Ch}(f_{3,n})-\mathrm{Ch}(f_{4,n})\neq 0$. That shows:
Proposition 4.12.
The metric measure space $(\mathbb{R}^{d},\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
This proves Theorem 2.
Remark 4.13.
For $d\geq 4$, the above example shows that we can find a Riemannian metric $g$ on $\mathbb{R}^{d}$ of the Geroch-Traschen class that induces a non-infinitesimally Hilbertian metric measure space.
Remark 4.14.
For a manifold with a $W^{1,p}_{\rm loc}$-Riemannian metric such that $p>d$, the Sobolev embedding theorem together with the results from [28] immediately give us that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian. Theorem 2 shows that $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ for $p<d-1$ is not sufficient for infinitesimal Hilbertianity. This works towards answering Question 3 leaving only the interval of $p\in[d-1,d]$, where it is unknown whether $g,g^{-1}\in L^{\infty}_{\rm loc}\cap W^{1,p}_{\rm loc}$ is sufficient to get infinitesimal Hilbertianity.
Remark 4.15.
One motivation behind the definition of infinitesimal Hilbertianity was to distinguish Finslerian structures from Riemannian ones, in particular in the context of synthetic lower Ricci curvature bounds. The distance in this example arises from a Riemannian metric of low regularity; however the corresponding metric measure space does not satisfy infinitesimal Hilbertianity. The directional dependence of the metric speed of curves established in (4.94) reveals the asymptotic shape of metric balls in the manifold and it is readily seen that the space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not locally Minkowskian (see [30] and [27] for a discussion of infinitesimal Hilbertianity for locally Minkowskian spaces). Moreover, the construction immediately implies that this space does not satisfy neither distributional lower Ricci curvature bounds nor the $\mathsf{CD}(K,\infty)$-condition for any $K\in\mathbb{R}$.
Statements and Declarations
Conflict of interest statement. The author has no relevant financial or non-financial interests to disclose. The author has no conflict of interest to declare that are relevant to the content of this article.
Data Availability. Data sharing is not applicable to this article as no new data were created or analysed in this study.